### v2

# Helfrich cylinders – instabilities, bifurcations and amplitude equations Alexander Meiners, Hannes Uecker

Abstract
Combining local bifurcation analysis with numerical continuation and bifurcation methods we study bifurcations from cylindrical vesicles described by the Helfrich equation with volume and area constraints, with a prescribed periodicity along the cylindrical axis. The bifurcating solutions are in two main classes, axisymmetric (pearling), and non–axisymmetric (coiling, buckling, and wrinkling), and depending on the spontaneous curvature and the prescribed periodicity along the cylinder axis we obtain different stabilities of the bifurcating branches, and different secondary bifurcations.
Keywords. Spontaneous curvature, Helfrich energy, bifurcation analysis, numerical continuation
MSC 35B32, 58J55, 65P30
1 Introduction
The (dimensionless) Canham–Helfrich model [Can70, Hel73] (or spontaneous curvature model) for vesicles with lipid bilayer membranes is based on the energy
$E(X)=\int_{X}(H-c_{0})^{2}\,{\rm d}S,$ (1)
where $X$ is a (possibly time $t$ dependent) 2D manifold embedded in ${\mathbb{R}}^{3}$ with mean curvature $H$, and where the parameter $c_{0}\in{\mathbb{R}}$ models an energetically favorable spontaneous (or preferred) curvature; it depends on the properties of the lipid bilayer membrane and can be zero, or positive, or negative. Roughly speaking, the aim is to minimize $E$ for prescribed area ${\cal A}(X)$ and enclosed volume ${\cal V}(X)$, i.e., under the constraints
 $\text{ $q_{1}(X):={\cal A}(X)-{\cal A}_{0}=0$ and $q_{2}(X):={\cal V}(X)-{\cal V}_{0}=0$}.$ (2)
There are numerous works on this, for instance using the minimization of $E$ to explain the shape of red blood cells, see, e.g., [SZ89, SBL90, Pet83, SL95, Sei97]. [^1] For $c_{0}=0$, $E$ is also called Willmore energy. More generally, a term $b\int_{X}K\,{\rm d}S$ can be added to $E$, where the parameter $b\geq 0$ is called saddle–splay modulus, but due to the Gauss–Bonnet theorem $\int KdS$ is a topological invariant for closed $X$, and hence can be dropped from the minimization of $E$, and the same holds for the cylindrical topology considered here. The minimization proceeds by introducing the Lagrangian
$F(X,\Lambda)=E(X)+\lambda_{1}({\cal A}(X)-{\cal A}_{0})+\lambda_{2}({\cal V}(X)-{\cal V}_{0}),$ (3)
with Lagrange multipliers $\Lambda=(\lambda_{1},\lambda_{2})\in{\mathbb{R}}^{2}$, where $\lambda_{2}$ is the pressure difference between inside and outside; since we choose the inner normal, here negative $\lambda_{2}$ corresponds to lower pressure inside; $\lambda_{1}$ corresponds to a surface tension, and can also have either sign; see, e.g., [BHAVMC17] for discussion.
From (3) we obtain the Euler–Lagrange equation
$G(X,\Lambda):=\Delta H+2H(H^{2}-K)+2c_{0}K-2(c_{0}^{2}+\lambda_{1})H-\lambda_{2}=0$ (4)
under variations in normal direction together with the constraints (2), where $K$ is the Gaussian curvature, and $\Delta$ the Laplace–Beltrami operator on $X$. Many works also use (1) to study instabilities of tubular vesicles, e.g., [ZcH89, BZM94, FTGB+99, TSA+01], in particular the so–called pearling and coiling instabilities, see Fig.1(a,b), and Remark 1 for a review of the (physics) literature.
(a) (b)      (c)
Refer to caption Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption
[…]
In, e.g., [NY12, RSS24], the constrained $L^{2}$–gradient flow associated to (1) is studied, which can be written as
$\begin{pmatrix}V_{t}(t)\\ 0\end{pmatrix}=\begin{pmatrix}G(X(t),\Lambda(t))\\ q(X(t))\end{pmatrix}$, (5)
where $q=(q_{1},q_{2})$ and $V_{t}(t)$ is the normal velocity of $X$, and with now time dependent Lagrange multipliers $\Lambda(t)$ which are determined from (2), such that
$\frac{\rm d}{{\rm d}t}E(X(t),\Lambda(t))=-\|V(t)\|_{L^{2}(X(t))}^{2},\quad\frac{\rm d}{{\rm d}t}{\cal A}(X(t))=0,\quad\text{and }\frac{\rm d}{{\rm d}t}{\cal V}(X(t))=0,$ (6)
as long as the flow exists. We also refer to [Gar13] and [BDGP23] for introductions to, reviews of, and extended bibliographies concerning geometric flows, both analytically and numerically, and also including mean–field approaches.
A cylinder ${\cal C}(r)$ of radius $r$ is a solution to (4) iff
$\lambda_{1}=\frac{1}{2}\left(\frac{1}{2r^{2}}-2r\lambda_{2}-2c_{0}^{2}\right).$ (7)
However, $E,{\cal A}$ and $\cal V$} are not well defined for an infinite cylinder, and in the following we restrict to cylinders ${\cal C}_{L}(r)$ of lengths $L$. For these we have the scaling $E({\cal C}_{L}(\alpha r);c_{0})=E({\cal C}_{\alpha L}(r);c_{0}/\alpha)$ and thus it is sufficient to consider $r=1$ with a variable length $L$, and we abbreviate ${\cal C}_{L}={\cal C}_{L}(1)$. After showing well–posedness of (5) for initial conditions $X_{0}$ close to ${\cal C}_{L}$, where we assume periodic boundary conditions along the cylindrical axis, we study bifurcation diagrams of steady states of (5). In Fig.1(c) we preview four types of solutions from branches bifurcating from cylinders, obtained from numerical continuation and bifurcation, namely again pearling (A) and coiling (B) like in (a) and (b), and buckling (C), and wrinkling (D), which break the symmetry of the cylinder in different ways.  [^2] While pearling is a standard name for A, the nomenclature for B and in particular C varies. For instance, B is sometimes called writhing or twisting, and in [NSS15], C is called wrinkling. We could not find explicit names for states like D in the “vesicles literature”, and our choice of “wrinkling” is inspired by [FVKG22], which treats a related model for arteries with a restoring force.  [^3] The symmetry group of the problem at the cylinder is $O(2)\times O(2)$, translations and reflections in $x$, and in $\varphi$, where $x$ is along the cylindrical axis and $\varphi$ the polar angle orthogonal to it. Additionally there are the trivial symmetries of translation in $y$ and $z$, the euclidian coordinates orthogonal to $x$, but these are mostly relevant from a numerical point of view, i.e., must be dealt with via phase conditions in the numerical continuation. The starting point for the bifurcation analysis is the dispersion relation $\mu=\mu(m,n)$ for the linearization of (5) at ${\cal C}_{L}$, which can be computed explicitly from an ansatz
$X={\cal C}_{L}+{\rm e}^{\mu t+{\rm i}(k_{m}x+n\varphi)}N$,   $m,n\in{\mathbb{Z}}$,    $k_{m}=2\pi m/L$.
We study these dispersion relations with $c_{0}$ and $\lambda_{2}$ (and $\lambda_{1}$ given by (7)) as primary parameters, and for suitable $c_{0}$ and $L$ we find (typically narrow) ranges in $\lambda_{2}$ of stability of the “trivial branch” $X={\cal C}_{L}$, with branch points at the ends. At one end, axisymmetric branches of pearling type A bifurcate, while at the other end the loss of stability is to B,C or D, depending on the parameters.
After identifying these instabilities, based on the well–posedness results for the flow we use a centre manifold reduction to derive amplitude equations (AEs) for the critical modes, and from these also obtain approximations of the bifurcating steady state branches. The derivation of these AEs exploits the $O(2)\times O(2)$ symmetry of the problem, but is rather non–standard due to the constraints (2), and in some sense our methods could be described as bifurcation without parameters [Lie15], see Remark 1.
Finally, we use the package pde2path [Uec21, MU24b] to numerically continue the primary bifurcating branches away from their bifurcation points from ${\cal C}_{L}$. Close to bifurcation these numerical results agree well with the AE description, and additionally we obtain the behavior and stability at larger amplitude, and secondary bifurcations. In summary, this yields rather complicated bifurcation diagrams (BDs), which we discuss in detail for three exemplary cases, namely $c_{0}=0,c_{0}=-1$, and $c_{0}=0.48$, with strong qualitative differences, and we complement these steady state BDs with numerical flows for (5) to give further indication of the typical behavior of (5).
Remark 1.1.
We start the linear stability analysis and also the numerical bifurcation analysis by continuation of the trivial solution (branch) ${\cal C}_{L}$ in the Lagrange multiplier $\lambda_{2}$ (with $\lambda_{1}$ given by (7)). We then also often continue the nontrivial branches (e.g., of type A–D from Fig.1) in $\lambda_{2}$ as a primary parameter, with fixed $\cal A$} but without fixing $\cal V$}. However, the stability is always computed wrt to (5), i.e., for the area and volume conserving flow. This point of view of allowing the volume $\cal V$} to change along the branch while considering stability under $\cal V$} conservation is motivated by the closed vesicle situation discussed for example in [JSL93] and [MU24b]. Roughly speaking, for closed vesicles, the sphere is the unique maximum of volume at fixed area. Hence volume and area preserving branches bifurcating from the sphere do not exist. The same holds for the cylinder if we also fix the length $L$. Cylinders can be destabilized preserving $\cal V$} and $\cal A$} when allowing $L$ to change, see, e.g., [ZcH89], but we found our method with fixed $L$ more convenient. $\rfloor$
Remark 1.2.
a) Following [BZM94], [GNPS96] considers the dynamics of pearling based on the Helfrich energy and treating the hydrodynamics in the tube via a lubrication approximation. This leads to a 6th order Cahn–Hilliard type equation for axially symmetric perturbations of the cylinder, with $\cal V$} conservation, and the critical wave numbers and growth rates of the pearling (including the propagation speed of pearling fronts) are carefully compared to [BZM94]. The coiling instability and the dynamics of coiling are discussed [FTGB+99, TGBKS01], and compared to the dynamics of pearling in [TSA+01]. For both, it is argued that polymers anchoring at lipid bilayers locally generate nonzero values for the spontaneous curvature $c_{0}$, and generate significant differences in effective area between inside and outside of the bilayer(s), which points at area difference elasticity (ADE) terms. In particular, in [TSA+01], measurements of diffusion/propagation speeds suggest that for local injection of polymers in the outer fluid, initially ADE terms (with bulk diffusion of polymers) give the right model, and spontaneous curvature terms (with surface diffusion of polymers) when reaching equilibrium, both with $\cal A$} and $\cal V$} conservation. On the other hand, [CR98] argues that in certain experiments $c_{0}\neq 0$ is a global quantity. Finally, an important feature of coiling is that it rather occurs for multilamellar tubes, and that in equilibrium the coiling is usually “tight”, as in Fig.1(b), and “stretched coils” as B in (c) only occur transiently. See also [Sta01] for an overview of these results.
Further experiments on coiling are discussed in [PQAEM09], with various hypothesis for the underlying mechanism, and [HEJ+14] discusses the role of the actin cortex (not considered in (5)) in suppressing pearling in actual cells. In [NSS15], tubular vesicles modeled by (1) with $c_{0}=0$ and $\cal A$} and $\cal V$} conservation, are subjected to external elongational and compressional flows, and numerical simulations are compared to experimental results. Depending on flow parameters (condensed to the capillary number), and the initial vesicle shape (including, e.g., the shape of the spherical caps at the ends of finite cylinders), pearling is observed for elongational flows, and buckling and coiling (called wrinkling in [NSS15]) for compressional flows. This is also complemented by linear stability analysis for infinite cylinders.
b) In our model (5) we do not consider any interaction with surrounding fluids; we treat the idealized infinitely long cylinders, or the idealization of a fixed spatial periodicity; and $c_{0}$ and the surface tension $\lambda_{1}$ and pressure difference $\lambda_{2}$ are always global. Hence, mismatches to experiments (small range of stability for periodic pearling, long wave nature of the buckling and coiling, see below) are not completely surprising, but we believe/hope that our analytical (instabilities of ${\cal C}_{L}$, derivation of the AEs) and numerical (continuation of the bifurcating branches to larger amplitude) discussion of (5) will also help the further modeling. $\rfloor$
Outline, and overview of results
In §2 we recall the needed differential geometry and rewrite (5) as an equation in the normal displacement $u$ from the cylinder. This gives a constrained quasilinear parabolic problem in $u$, for which we use maximal regularity in Sobolev spaces to obtain local existence close to ${\cal C}_{L}$. From this we also obtain that nonlinear stability follows from linearized stability, and since ${\cal C}_{L}$ has constant mean and Gauss curvatures, the eigenvalues and eigenfunctions of the linearized problem can be computed explicitly allowing an analytical treatment of the instabilities of ${\cal C}_{L}$, and of the bifurcating branches via amplitude equations in §3 and §4. This is straightforward, but cumbersome due to the three parameters $c_{0},L$ and $\lambda_{2}$, and requires a careful look at the area and volume constraints (2) already in the linearization. The results of the linear stability and AE analysis can be summarized as follows:
  1. 1.
    The coiling/buckling instability of ${\cal C}_{L}$ is always of long wave type, i.e., the primary coiling/buckling instability has the maximal period given by the length $L$. This contradicts many experimental results for coiling, where the instability seems of finite wavelength type, and after onset the cylinders coils up tightly as in Fig.1(b). We could not find clear experimental results on buckling of vesicles.
  2. 2.
    For some (positive) $c_{0}$ the primary pearling instability can be of finite wave number type, i.e., independent of $L$, corresponding to Fig.1(a). However, see 5. below.
  3. 3.
    The wrinkling instability (homogeneous along the cylindrical axis) naturally is independent of $L$.
  4. 4.
    Due to the constraints, also the derivation and interpretation of the AEs is non–standard. In particular, the pearling and wrinkling branches bifurcating at stability loss of ${\cal C}_{L}$ are always stable. For the coiling vs buckling at stability loss of ${\cal C}_{L}$, always exactly one of these bifurcating branches is stable, and the instability of the unstable one is wrt perturbations into direction of the stable one.
In §5 we complement our analysis with numerical continuation and bifurcation methods, and also with some numerical flows for (5), aka direct numerical simulation (DNS). As expected and already said, near bifurcation the numerics agree well with the analysis, but one remarkable result is the following:
  1. 5.
    In case of finite wave number pearling the stability range of the bifurcating pearling branch rather strongly depends on $L$: The pearling branch becomes unstable via exchange of stability with a bifurcating branch of a “localized single pearl”, and the associated bifurcation point moves closer to the primary bifurcation with increasing $L$. This suggests that finite wavelength $O(1)$ amplitude pearling as in Fig.1(a) does not arise via bifurcations at stability loss of ${\cal C}_{L}$ in (5).
See §5 for details of this, and also for secondary bifurcations from other primary branches. On the other hand:
  1. 6.
    For the case of “large” $c_{0}$ ($c_{0}>1/2$, where ${\cal C}_{L}$ is not stable for any $(\lambda_{1},\lambda_{2})$) we do find finite wavelength $O(1)$ pearling, but not via bifurcation from ${\cal C}_{L}$. Instead we use DNS to converge to such solutions, which we subsequently can continue again in parameters. However, we only present this as an outlook in §5.4, and leave a systematic study of this setting for future work.
Another interesting result from the numerical flows is that cylindrical shapes can be more stable than linearized stability analysis of steady solutions $(X,\lambda_{1},\lambda_{2})$ at fixed ${\cal A}_{0}$ and ${\cal V}_{0}$ suggests. Namely:
  1. 7.
    Non area and/or volume conserving perturbations, of, e.g., unstable coiling, may yield initial conditions which converge to nearby stable coiling at the new conserved $({\cal A}_{0},{\cal V}_{0})$ by mainly adjusting the surface tension $\lambda_{1}(t)$ and the pressure $\lambda_{2}(t)$, with rather negligible changes of $X(t)$.
In §6 we summarize our analysis again, and in the Appendix we collect some details of the derivation of the AEs, and some remarks on the numerical algorithms.
2 Preliminaries and well-posedness of the flow
We introduce the basic differential geometric preliminaries in order to write (5) as a constrained parabolic quasi-linear equation in the normal displacement $u$ of the cylinder. This construction is as in [LS16], and hence we refrain from extensive computational details and focus on the structure and mapping properties of the resulting equations. From these we obtain maximal regularity results and hence local existence for (5). For general background on differential geometry we recommend [Tap16, UY17].
2.1 Preliminaries
Let ${\cal C}(r)$ be the two dimensional unbounded cylinder embedded in ${\mathbb{R}}^{3}$ of radius $r>0$ with the unbounded direction along the $x$ axis. ${\cal C}(r)$ can be considered as the embedding
$\begin{array}[]{rl}\Phi:{\mathbb{R}}\times\mathbb{T}&\rightarrow{\mathbb{R}}^{3}\\ (x,\varphi)&\mapsto\begin{pmatrix}x,r\cos(\varphi),r\sin(\varphi)\end{pmatrix}^{T}\end{array}$ (10)
of the product manifold ${\mathbb{R}}\times\mathbb{T}$, where $\mathbb{T}$ is the one dimensional topological torus with length $2\pi$ endowed with the metric $d_{\mathbb{T}}(\varphi,\vartheta)=\min\{|\varphi-\vartheta|,2\pi-|\varphi-\vartheta|\}$ for all $\varphi,\vartheta\in\mathbb{T}$. Then as the pullback metric of the ${\mathbb{R}}^{3}$ to ${\cal C}(r)$ we have $g_{0}=\,{\rm d}x^{2}+r^{2}\,{\rm d}\varphi^{2}$. The normal vector (field) $N$ of ${\cal C}(r)$ is oriented inwards, wrt $\Phi$, hence $H({\cal C}_{r})=\frac{1}{2r}$.
For a function $u:{\cal C}(r)\rightarrow{\mathbb{R}}$, for the remainder of this subsection assumed as smooth as necessary, we consider surfaces
$X={\cal C}(r)-u\,N,$
assuming $u<r$, such that $X$ is still an embedding of ${\mathbb{R}}\times\mathbb{T}$; explicitly,
$\tilde{\Phi}(x,\varphi)=\Phi(x,\varphi)+\tilde{u}(x,\varphi)(0,\cos(\varphi),\sin(\varphi))^{T},$
where $\tilde{u}=u\circ\Phi$. This gives the metric
$(g_{ij})=\begin{pmatrix}1+u_{x}^{2}&u_{x}u_{\varphi}\\ u_{x}u_{\varphi}&(r+u)^{2}+u^{2}_{\varphi}\end{pmatrix}\quad\text{with inverse }(g^{ij})=\frac{1}{g}\begin{pmatrix}(r+u)^{2}+u^{2}_{\varphi}&-u_{x}u_{\varphi}\\ -u_{x}u_{\varphi}&1+u_{x}^{2}\end{pmatrix}$
with $g=\det(g_{ij})=(r+u)^{2}(1+u_{x}^{2})+u_{\varphi}^{2}$. In the same explicit fashion one can calculate the second fundamental form
$(h_{ij})=\frac{1}{\sqrt{g}}\begin{pmatrix}(r+u)u_{xx}&(r+u)u_{x\varphi}-u_{x}u_{\varphi}\\ (r+u)u_{x\varphi}-u_{x}u_{\varphi}&(r+u)(u_{\varphi\varphi}-(r+u))-2u_{\varphi}^{2}\end{pmatrix}$,
and this gives explicit formulas for the mean and Gaussian curvatures,
$H(u)$ $=\frac{1}{2g^{3/2}}\left(u_{\varphi}^{2}-(r+u)\big{[}((r+u)^{2}+u_{\varphi}^{2})u_{xx}+(1+u_{x}^{2})u_{\varphi\varphi}-2u_{x}u_{\varphi}u_{x\varphi}\big{]}\right)+\frac{1}{2g^{1/2}},$ (11)
$K(u)$ $=\frac{1}{g^{2}}\left((r+u)u_{xx}\Big{[}(r+u)\big{(}u_{\varphi\varphi}-(r+u)\big{)}-2u_{\varphi}^{2}\Big{]}-\big{(}(r+u)u_{x\varphi}-u_{x}u_{\varphi}\big{)}^{2}\right).$ (12)
We denote by $\Delta_{g}$ the negative definite Laplace–Beltrami operator
$\Delta_{u}=\frac{1}{g^{1/2}}\left({\partial}_{x}\bigg{[}\frac{(r+u)^{2}+u_{\varphi}^{2}}{g^{1/2}}{\partial}_{x}-\frac{u_{x}u_{\varphi}}{g^{1/2}}{\partial}_{\varphi}\bigg{]}+{\partial}_{\varphi}\bigg{[}-\frac{u_{x}u_{\varphi}}{g^{1/2}}{\partial}_{x}+\frac{1+u_{x}^{2}}{g^{1/2}}{\partial}_{\varphi}\bigg{]}\right),$ (13)
which is self adjoint in $L^{2}(X)$. Furthermore, the normal velocity is
$V_{t}=-\left\langle\frac{\rm d}{{\rm d}t}X,N\right\rangle=\frac{(r+u)u_{t}}{g^{1/2}}=\eta(u)u_{t}$ (14)
where $\eta(u)>0$ for $u<r$, as we assumed. By all of this we can express (5) as a PDE in $u$, and with a slight abuse of notation we rewrite (5) as
$\eta(u){\partial}_{t}u$ $=G(u,\Lambda),$ (15a)
$0$ $=q(u),$ (15b)
where $G$ has the quasi-linear structure $G(u,\Lambda){=}B(u)u+C(u)$, where $B(u){=}\sum_{|\alpha|=4}b_{\alpha}{\partial}_{x}^{\alpha_{1}}{\partial}_{\varphi}^{\alpha_{2}}$, where $b_{\alpha}=b_{\alpha}(u,\nabla u,\nabla^{2}u)$, see (27), and where similarly $C$ is a nonlinear operator of order up to three, but linear in the third derivatives.
Naturally, we shall need linearizations of (15) at ${\cal C}_{L}(r)$ and at $X(u)={\cal C}_{L}(r)-uN$. For this we first note the formulas for derivatives of area $\cal A$} and volume $\cal V$}, where we distinguish between the derivative of a functional ${\cal H}(u)$ acting on $v$ in the $L^{2}$ scalar product, denoted by $\left\langle{\partial}_{u}{\cal H},v\right\rangle$, and the integral kernel ${\partial}_{u}{\cal H}$ as a differential operator on $X$.
Lemma 2.1.
For $u,v\in C^{\infty}_{\text{per}}({\cal C}_{L}(r))$ sufficiently small, the area of $X(u)={\cal C}_{L}(r)-u\,N$ expands as ${\cal A}(u+v)={\cal A}(u)+\left\langle{\partial}_{u}{\cal A}(u),v\right\rangle+\frac{1}{2}\left\langle{\partial}_{u}^{2}{\cal A}(u)v,v\right\rangle+O(\|v\|^{3})$ with
$\left\langle{\partial}_{u}{\cal A}(u),v\right\rangle{=}2\int_{X}H(u)\eta(u)v\,{\rm d}S,\ \ \left\langle{\partial}_{u}^{2}{\cal A}(u)v,v\right\rangle{=}\int_{X}-(\Delta_{u}v+2(H^{2}(u)-K(u))v)v\,{\rm d}S.$ (16)
The volume
${\cal V}(u)=\frac{1}{2}\int_{X}\frac{(r+u)^{2}}{g^{1/2}(u)}\,{\rm d}S$ (17)
expands as ${\cal V}(u+v)={\cal V}(u)+\left\langle{\partial}_{u}{\cal V}(u),v\right\rangle+\frac{1}{2}\left\langle{\partial}_{u}^{2}{\cal V}(u)v,v\right\rangle+O(\|v\|^{3})$ with
$\left\langle{\partial}_{u}{\cal V}(u),v\right\rangle=\int_{X}\eta(u)v\,{\rm d}S,\qquad\left\langle{\partial}_{u}^{2}{\cal V}(u)v,v\right\rangle=\int_{X}g^{-1/2}(u)v^{2}\,{\rm d}S.$ (18)
Remark 2.2.
Since $\Lambda$ appears linearly in $G(u,\Lambda)$, the first derivative $\frac{\rm d}{{\rm d}t}q(u)=0$ gives a linear system of equations
$\mathcal{D}\Lambda=\begin{pmatrix}\left\langle{\partial}_{u}E,{\partial}_{u}{\cal A}(u)\right\rangle\\ \left\langle{\partial}_{u}E,{\partial}_{u}{\cal V}(u)\right\rangle\end{pmatrix}$, (19)
for $\Lambda$, where
$\mathcal{D}:=\left\langle{\partial}_{u}q(u),{\partial}_{\Lambda}G(u,\lambda)\right\rangle=\begin{pmatrix}4\int_{X}H^{2}(u)\eta^{2}(u)\,{\rm d}S&2\int_{X}H\eta^{2}(u)\,{\rm d}S\\ 2\int_{X}H\eta^{2}(u)\,{\rm d}S&\int_{X}\eta^{2}(u)\,{\rm d}S\end{pmatrix}$. (20)
If $\det\mathcal{D}=4\left(\int_{X}\eta^{2}\,{\rm d}S\int_{X}H^{2}\eta^{2}\,{\rm d}S-\left(\int_{X}H\eta^{2}\,{\rm d}S\right)^{2}\right)\neq 0$, then the implicit function theorem gives that in a neighborhood of $u$ there is a unique solution $\Lambda=\Lambda(u)$. Following [NY12], inserting $\Lambda(u)$ in (15a) gives a parabolic integro differential equation. However, if $u=0$, then $\det{\cal D}=0$, but the right hand side of (19) is orthogonal to $\ker{\cal D}$ (more generally, this holds at any constant mean curvature surface). Then a solution $\Lambda$ to (19) exists, and $\left\langle\Lambda_{\cal C},\ker{\cal D}^{T}\right\rangle=0$ gives that $2\lambda_{1}=\lambda_{2}$ and $\lambda_{1}=\frac{1}{3{\cal A}_{0}}\left\langle{\partial}_{u}E,1\right\rangle$, i.e., $\ker{\cal D}$ corresponds to the family of cylinders in (7); see [NY12] for further discussion (for the case of closed vesicles of spherical topology).
Here we shall proceed similarly for the cylindrical vesicles; we start with well–posedness of (15a) for fixed $\Lambda\in{\mathbb{R}}^{2}$, i.e., with inactive constraints (15b). Then we first switch on one constraint, wlog $q_{1}(u)=0$ and free $\lambda_{1}=\lambda_{1}(t)$, and subsequently consider the full constraints $q(u)=(q_{1}(u),q_{2}(u))=0$ with free $\Lambda(t)=(\lambda_{1}(t),\lambda_{2}(t))$. $\rfloor$
2.2 Well-posedness
Given an initial condition $u_{0}\in{\cal X}\subset L^{2}({\cal C}_{L}(r))$, i.e., $u_{0}=u_{0}(x,\varphi)$ with periodic boundary conditions in $x$, where ${\cal X}=H^{m}_{\text{per}}({\cal C}_{L}(r))$ with the choice $m\in(3,4)$ explained below, we search for solutions $(t,x,\varphi))\mapsto(u(t,x,\varphi),\Lambda(t))$ of (15) on $0,T)\times{\cal C}_{r}(L)$. Here ${\cal C}_{L}(r)$ is topologically equivalent to a torus, and the periodic Sobolev spaces of order $m\geq 0$ are
$H^{m}_{\text{per}}=H^{m}_{\text{per}}({\cal C}_{L}(r))=\{v\in L^{2}({\cal C}_{L}(r)):\sum_{k\in\frac{2\pi}{L}{\mathbb{Z}}\times{\mathbb{Z}}}(1+|k|^{2})^{m}|\hat{v}_{k}|^{2}<\infty\},$ (21)
where $\hat{v}_{k}=\left\langle v,e^{ik\cdot(x,\varphi)}\right\rangle$. Clearly, $H^{m_{1}}_{\text{per}}{\subset}H^{m_{2}}_{\text{per}}$ for $m_{1}{\geq}m_{2}$, and $H^{m_{1}}_{\text{per}},H^{m_{2}}_{\text{per}}]_{\theta}{=}H^{(1-\theta)m_{1}+\theta m_{2}}_{\text{per}}$ for the associated real interpolation spaces, see [LM72] and [Tri78, §2.4.2], and $\nabla\in{\cal L}(H^{m}_{\text{per}},H^{m-1}_{\text{per}})$.
In the following, for $m>1$ we restrict the Sobolev spaces to open subsets
$V^{m}_{\varepsilon}:=H^{m}_{\text{per}}({\cal C}_{L}(r))\cap\{|u|<r-\varepsilon\}$ (22)
with some $0<\varepsilon<r$. Then ${\cal C}_{L}(r)-uN$ is an embedding of ${\cal C}_{L}(r)$, and we can easily state local Lipschitz properties of the operators involved in (15). For convenience we recall some Sobolev embedding results, which hold for compact manifolds with boundary as for compact domains in ${\mathbb{R}}^{n}$, see, e.g., [Aub82]. [^4] The situation of non compact manifolds is more complicated, but if the manifold is complete, has a positive injection radius and bounded sectorial curvature, then the same results holds as well. In fact, in the following analysis near ${\cal C}_{L}(r)$ we do not strictly need the manifold versions of Sobolev embeddings due to the global coordinates (10) also underlying (21), and hence we basically state the more general embedding results for possible generalization to other manifolds. In 2D, $H^{m}(X)\hookrightarrow L^{\infty}(X)$ for $m>1$ embeds continuously, while $H^{1}(X)$ does not. Also, for $m=1+\alpha$, $H^{m}(X)\hookrightarrow C^{0,\alpha}(X)$, hence the point evaluation $u(x,\phi)$ and in particular the sets $\{|u(x,\varphi)|<r-\varepsilon\}$ are well defined for $m>1$.
2.2.1 Well-posedness of the PDE part
Following Remark 2.1 we first consider the PDE part (15a) with fixed $\Lambda\in{\mathbb{R}}^{2}$, i.e.,
$\eta(u){\partial}_{t}u=B(u)u+C(u),$ (23)
where $B(u)$ collects the 4th order derivatives $b_{\alpha_{1},\alpha_{2}}(u){\partial}_{x}^{\alpha_{1}}{\partial}_{\phi}^{\alpha_{2}}$, $|\alpha|:=\alpha_{1}+\alpha_{2}=4$, namely
$\begin{array}[]{l}b_{4,0}(u)=g^{-2}((r{-}u)^{2}+u_{\phi}^{2})^{2},\quad b_{3,1}(u)=-g^{-2}(4u_{x}u_{\phi}((r+u)^{2}+u_{\phi}^{2}),\\ b_{2,2}(u)=g^{-2}(2((r{-}u)^{2}+u_{\phi}^{2})(1+u_{\phi}^{2})+4u_{x}^{2}u_{\phi}^{2},\\ b_{1,3}(u)=-4g^{-2}(u_{x}u_{\phi}(1+u_{x}^{2}),\quad b_{0,4}=g^{-2}(1+u_{x}^{2})^{2},\end{array}$ (27)
cf.[LS16], where our $B(u)$ appears as the principal part of the 4th order operator $P(u)$ in [LS16].
Theorem 2.3.
Let $m\in(3,4)$ and fix $\Lambda\in{\mathbb{R}}^{2}$. Then for all $u_{0}\in V^{m}_{\varepsilon}$ there exists a $t_{0}=t_{0}(u_{0})$ and a unique solution $u(\cdot,u_{0})\in L^{2}((0,t_{0}),H^{4}(\mathcal{C}))\cap H^{1}((0,t_{0}),L^{2}(\mathcal{C}))\cap C([0,t_{0}],V_{\varepsilon}^{m}))$ of (23).
To prove Theorem 2.3 we apply [MRS13, Theorem 3.1], which treats abstract quasilinear parabolic problems of the form
${\partial}_{t}u=A(u)u+F(u),\quad t>0,\quad u(0)=u_{0}\in{\cal X},$ (28)
with Gelfand triple $X_{0}\hookrightarrow{\cal X}\hookrightarrow X_{1}=D(A(u))$, i.e., phase space $\cal X$}. For this we could rewrite (23) as ${\partial}_{t}u{=}\widetilde{B}(u)u{+}\widetilde{C}(u)$ with $\widetilde{B}(u){=}\eta(u)^{-1}B(u)$ and $\widetilde{C}(u){=}\eta(u)^{-1}C(u)$, but since $\eta(u)=(1+u)g^{-1/2}$ contains at most first derivatives it will be clear from the following that we can as well keep $\eta(u)$ in (23) on the lhs and consider the rhs of (23) in the setting of (28). As already indicated above, we choose the base space $X_{0}=L^{2}$ and hence $X_{1}=D(B(u))=H^{4}$, and we need to find an open subset $V=V^{m}_{\varepsilon}$ of the phase space $\cal X$} with $X_{0}\subset{\cal X}\subset X_{1}$ such that
  • (W1)
    $B:V\to L(X_{1},X_{0})$ and $C:V\to X_{0}$ are locally Lipschitz;
  • (W2)
    for all $u\in V$, $B(u)$ generates a $C_{0}$ (strongly continuous) analytic semigroup in $X_{0}$.
By [MRS13, Theorem 3.1], (23) is then locally well posed in $V$, i.e., there exists a $t_{0}>0$ and a unique solution $u\in C([0,t_{0}),V)\cap H^{1}((0,t_{0}),X_{0})\cap L^{2}((0,t_{0}),H^{4})$. To show (W1) we use Sobolev embeddings and the structures of $B$ and $C$, while (W2) essentially follows from $B(u)$ being uniformly elliptic (in the 4th order sense). In detail, the symbol $\sigma[B(u)](x,\phi,\xi)$ satisfies
$\sigma[B(u)](x,\phi,\xi)\geq g^{-2}((r-u)^{2}\xi_{1}^{2}+\xi_{2}^{2})^{2},$ (29)
and hence for $\|u\|_{\infty}<r-\varepsilon$, there exists constants $c_{1},c_{2}$ such that $c_{1}\leq\sigma[B(u)](x,\phi,\xi)\leq c_{2}$ for all $(x,\phi,\xi)\in{\cal C}_{L}\times{\mathbb{R}}^{2}$ with $\|\xi\|=1$.
Lemma 2.4.
For $m\in(3,4)$, $B(\cdot)\in C^{\infty}(V^{m}_{\varepsilon},{\cal L}(H^{4}_{\text{per}},L^{2}))$ and $C\in C^{\infty}(V^{m}_{\varepsilon},L^{2})$. In addition, the rhs of (15) is an analytic map from $V^{4}_{\varepsilon}\times{\mathbb{R}}^{2}$ into $L^{2}$.
Proof.
From the Sobolev embeddings, a function $F(u)=u^{j}$ is well defined for any $j$ on $H^{m}$ if $m>1$, and the map $u\mapsto g(u)$ is smooth from $H^{m}(X)$ to $H^{m-1}(X)$ for $m>2$, and on $V^{m}$ bounded away from zero. Hence $g^{-s}\in L^{\infty}(X)$ for $s>0$ and due to the embedding $L^{\infty}(X)\hookrightarrow L^{2}(X)$, these terms are bounded in $L^{2}$. For $u$ small enough $g^{-s}$ can be expanded into a convergent power series, which gives the analyticity.
The coefficients (27) of $B(u)$ are well defined for $u\in V^{m}_{\varepsilon}$ with $m>3$ since $u,{\partial}_{x}u,{\partial}_{\phi}u\in L^{\infty}$. Actually, $m>2$ is enough here, and the condition $m>3$ comes from $C(u)$, which has summands $c_{\alpha}(u,\nabla u,\nabla^{2}u)=\tilde{c}_{\alpha}\,u^{j_{1}}({\partial}_{x}u)^{j_{2}}({\partial}_{\varphi}u)^{j_{3}}({\partial}_{x}^{2}u)^{j_{4}}({\partial}_{\varphi}^{2}u)^{j_{5}}({\partial}_{x}{\partial}_{\varphi}u)^{j_{6}}g^{-s}(u,\nabla u),$ (nonlinear terms up to second order), and $d_{\alpha}(u,\nabla u){\partial}_{x,\phi}^{\alpha}u,|\alpha|=3,$ where $d_{\alpha}$ is nonlinear but the third order terms (from $\Delta H(u)$ with $H(u)$ from (11)) only appear linearly. The maps $b_{\alpha}(u,y)$ are analytic on $\{|u|<r-\varepsilon\}\times{\mathbb{R}}^{2}$, and so are the $c_{\alpha}(u,y,z)$ on $\{|u|<r-\varepsilon\}\times{\mathbb{R}}^{2}\times{\mathbb{R}}^{4}$, and the $d_{\alpha}$ on $\{|u|<r-\varepsilon\}\times{\mathbb{R}}^{2}$. By introducing a perturbation $u+\delta\,v$ for $v\in H^{\alpha}$, and computing the derivative wrt $\delta$, we have, e.g.,
${\partial}_{u}c_{\alpha}(u,\nabla u,\nabla^{2}u)=({\partial}_{x}c_{\alpha}(x,y,z)v+{\partial}_{y}c_{\alpha}(x,y,z)\cdot\nabla v+{\partial}_{z}c_{\alpha}(x,y,z)\cdot\nabla^{2}v)|_{(x,y,z)=(u,\nabla u,\nabla^{2}u)}.$
The structure of ${\partial}_{u}c_{\alpha}$ is hence the same as that of $c_{\alpha}$, and similar for the $b_{\alpha}$ and $d_{\alpha}$ such that these ”coefficients” respectively summands are well defined and finite in $L^{\infty}$. This can be iterated for higher derivatives, to find the same result, and, e.g., for $w\in H^{4}$ and $u\in V^{m}_{\varepsilon}$,
$\|{\partial}_{u}B(u)w\|_{L^{2}}\leq C\|w\|_{H^{4}}.$
This gives the first two assertions, and setting $u=w$ gives the analyticity of $G(u,\Lambda)$, where the analyticity with respect to $\Lambda$ is clear since $\Lambda$ appears linearly. ∎
To show (W2) we use standard theory for sectorial operators [Hen81, Lun95]. Let $u\in V^{m}_{\varepsilon}$. Since $D(B(u))=H^{4}$ is dense in $X=L^{2}$, it is sufficient to show that $B(u)$ is sectorial in $L^{2}$, i.e., that the spectrum of $B(u)$ is contained in a sector $S_{\alpha,\omega}$ in $\mathbb{C}$} with tip at some $\omega\in{\mathbb{R}}$ of angle $\alpha<\pi$ opening to the left, and that there exists and $M>0$ such that the resolvent estimate
$\|(\lambda-B(u))^{-1}\|\leq\frac{M}{|\lambda-\omega|}$ (30)
holds for all $\lambda\not\in S_{\alpha,\omega}$. Essentially, this follows from the uniform ellipticity (29). In more detail, we show that $B(u)$ induces a coercive sesquilinear form on $H^{2}$, from which (W2) follows via the Lax–Milgram Lemma, see, e.g., [Sch24, Corollary 2.29].
Lemma 2.5.
For $u\in V^{m}_{\varepsilon}$ and $\lambda$ sufficiently large, the operator $(\lambda-B(u))$ induces a coercive sesquilinear form on $H^{2}_{\text{per}}({\cal C}_{L}(r))$, i.e., there exists $\omega,C$ independent of $u\in V^{m}_{\varepsilon}$ such that for $\lambda>\omega$ we have
$\int_{{\cal C}_{r}(L)}((\lambda w-B(u)w)w\,{\rm d}S\geq C\|w\|^{2}_{H^{2}_{\text{per}}({\cal C}_{r}(L))}.$ (31)
Proof.
The idea is to use (29), but since $B(u)$ is not in divergence form we need to apply the standard trick of writing, e.g.,
$b_{4,0}(u){\partial}_{x}^{4}w={\partial}_{x}^{2}(b_{4,0}(u){\partial}_{x}^{2}w)-{\partial}_{x}(\widetilde{b}_{4,0}(u){\partial}_{x}^{2}w),$
where $\widetilde{b}_{4,0}(u)={\partial}_{x}b_{4,0}(u)$ is still bounded in $L^{\infty}$, i.e., $\|\widetilde{b}_{4,0}(u)\|_{\infty}\leq C_{4,0}$. Therefore, using $|w_{xx}w_{x}|\leq\frac{\delta_{1}}{2}w_{xx}^{2}+\frac{1}{2\delta_{1}}w_{x}^{2}$, $\delta_{1}>0$, and similarly standard $L^{2}$–interpolation estimates such as $\int_{{\cal C}_{L}(r)}w_{x}^{2}\,{\rm d}S\leq\int_{{\cal C}_{L}(r)}\frac{\delta_{2}}{2}w_{xx}^{2}+\frac{1}{2\delta_{2}}w^{2}\,{\rm d}S$, we obtain, e.g.,
$\left\langle(\lambda-b_{4,0}(u))w,w\right\rangle$ $=\lambda\int w^{2}\,{\rm d}S+\int b_{4,0}(u)w_{xx}^{2}+\widetilde{b}_{4,0}(u)w_{xx}w_{x}\,{\rm d}S$
$\geq\left(c_{1}-\frac{1}{2}(\delta_{1}+\frac{\delta_{2}}{\delta_{1}})\right)\int w_{xx}^{2}\,{\rm d}S+\left(\lambda-\frac{C_{4,0}}{4\delta_{1}\delta_{2}}\right)\int w^{2}\,{\rm d}S,$
with $c_{1}>0$ from after (29). Proceeding similarly for the other terms from (27) and choosing suitable $\delta_{1,2}$ and then $\omega$ sufficiently large yields (31), uniformly in $u\in V^{m}_{\varepsilon}$. ∎
In summary we have verified (W1) and (W2) with $V=V^{m}_{\varepsilon}$ with $m\in(3,4)$ corresponding to $\vartheta\in(3/4,1)$ for the interpolation space $L^{2},H^{4}]_{\vartheta,2}=H^{m}$, and Theorem 2.3 now folows from [MRS13, Theorem 3.1].
2.2.2 The constrained flows
In order to take the constraints into account, we first consider only one active constraint namely $q_{1}(u)=0$, and fix $\lambda_{2}$, while the same arguments also apply for the case of active $q_{2}$ with fixed $\lambda_{1}$. Defining $\overline{H}=\frac{H}{\|H\|}_{L^{2}}$ there exists a $\sigma_{1}$ such that $(-{\partial}_{u}E+\lambda_{2}{\partial}_{u}{\cal V}+\lambda_{1}H)=-(-{\partial}_{u}E+\lambda_{2}{\partial}_{u}{\cal V}+\sigma_{1}\overline{H})$, and we define the projection ${\cal P}_{1}:L^{2}({\cal C}_{r}(L))\rightarrow\mathrm{span}\{{\partial}_{u}{\cal A}(u)\}^{\bot}$ by ${\cal P}_{1}f=f-\left\langle\overline{H},f\right\rangle\overline{H}$, and consider
$V_{t}={\cal P}_{1}({\partial}_{u}E-\lambda_{2}{\partial}_{u}{\cal V}).$ (32)
Then clearly $\frac{\rm d}{{\rm d}t}{\cal A}(u)=2\int_{{\cal C}_{r}(L)}H(u)V_{t}\,{\rm d}S=0$ and for $u\in V^{m}_{\varepsilon}$ we find
$\lambda_{1}$ $=\frac{1}{\int_{{\cal C}_{r}(L)}H^{2}\,{\rm d}S}\left\langle{\partial}_{u}E{-}\lambda_{2}{\partial}_{u}{\cal V},2H\right\rangle$
$=\frac{1}{\int_{{\cal C}_{r}(L)}H^{2}\,{\rm d}S}\int_{{\cal C}_{r}(L)}(\Delta H{+}2H(H^{2}{-}K){+}c_{0}K{-}2c_{0}H{-}\lambda_{2}{\partial}_{u}{\cal V})H\,{\rm d}S$
$=\frac{1}{\int_{{\cal C}_{r}(L)}H^{2}\,{\rm d}S}\int_{{\cal C}_{r}(L)}-|\nabla H|^{2}+2H^{2}(H^{2}{-}K)+c_{0}HK-2c_{0}H^{2}-\lambda_{2}{\partial}_{u}{\cal V}H\,{\rm d}S.$
Importantly, the term $|\nabla H|^{2}$ includes third order derivatives of $u$ but only quadratically with lower order “coefficients” and hence $\int|\nabla H(u)|^{2}\,{\rm d}S$ is well defined and analytic for $u\in H^{m}_{\text{per}}$ for $m>3$. This yields the analogous result as Theorem 2.3.
Theorem 2.6.
For $u_{0}\in V^{m}_{\varepsilon}$ there exists a $t_{0}=t_{0}(u_{0})$ and a unique solution
$(u(\cdot,u_{0}),\lambda_{1})\in(L^{2}((0,t_{0}),H^{4}(\mathcal{C}))\cap H^{1}((0,t_{0}),L^{2}(\mathcal{C}))\cap C([0,t_{0}),V_{\varepsilon}^{s})))\times C([0,t_{0}),{\mathbb{R}}^{1}).$
of $\eta(u){\partial}_{t}u=G(u,(\lambda_{1},\lambda_{2}))$ under the constraint $q_{1}(u)={\cal A}({\cal C}_{L}-uN)-{\cal A}_{0}=0$.
If $\det{\cal D}(u)\neq 0$, then we can consider the full constraints in (15) by projecting $V_{t}={\partial}_{u}E$ onto the area and volume preserving flow. We construct an orthonormal basis of $\mathrm{span}\{{\partial}_{u}{\cal A},{\partial}_{u}{\cal V}\}=\mathrm{span}\{H,(r+u)g^{-1/2}(u)\}$ by Gram-Schmidt as
$\overline{{\partial}_{u}{\cal V}}$ $=\frac{1}{\|{\partial}_{u}{\cal V}\|_{L^{2}}}{\partial}_{u}{\cal V}=\frac{1}{\|(r+u)g^{-1/2}(u)\|_{L^{2}}}(r+u)g^{-1/2}(u)$
$\overline{{\partial}_{u}{\cal A}}$ $=\frac{1}{\|H-\left\langle\overline{{\partial}_{u}{\cal V}},H\right\rangle\|_{L^{2}}}(H-\left\langle\overline{{\partial}_{u}{\cal V}},H\right\rangle)=\frac{1}{\sqrt{\|H\|_{L^{2}}^{2}-\left\langle\overline{{\partial}_{u}{\cal V}},H\right\rangle^{2}}}\left(H-\left\langle\overline{{\partial}_{u}{\cal V}},H\right\rangle\right).$
We define the projection
${\cal P}_{2}f=f-\left\langle\overline{{\partial}_{u}{\cal V}},f\right\rangle\overline{{\partial}_{u}{\cal V}}-\left\langle\overline{{\partial}_{u}{\cal A}},f\right\rangle\overline{{\partial}_{u}{\cal A}},$
and obtain $\frac{\rm d}{{\rm d}t}{\cal A}=\left\langle{\partial}_{u}{\cal A},V_{t}\right\rangle=0$ and $\frac{\rm d}{{\rm d}t}{\cal V}=\left\langle{\partial}_{u}{\cal V},V_{t}\right\rangle=0$ for the flow $V_{t}={\cal P}_{2}{\partial}_{u}E$. Since $\det{\cal D}(u)\neq 0$, (19) is uniquely solvable also in a small neighborhood of $u$. So by inverting ${\cal D}(u)$ the Lagrange multipliers are
$\lambda_{1}$ $=\frac{1}{\det{\cal D}}\big{(}\left\langle{\partial}_{u}{\cal V},{\partial}_{u}{\cal V}\right\rangle\left\langle{\partial}_{u}E,{\partial}_{u}{\cal A}\right\rangle-\left\langle{\partial}_{u}{\cal A},{\partial}_{u}{\cal V}\right\rangle\left\langle{\partial}_{u}E,{\partial}_{u}{\cal V}\right\rangle\big{)}$
$\lambda_{2}$ $=\frac{1}{\det{\cal D}}\big{(}\left\langle{\partial}_{u}{\cal V},{\partial}_{u}{\cal V}\right\rangle\left\langle{\partial}_{u}E,{\partial}_{u}{\cal V}\right\rangle-\left\langle{\partial}_{u}{\cal A},{\partial}_{u}{\cal V}\right\rangle\left\langle{\partial}_{u}E,{\partial}_{u}{\cal A}\right\rangle\big{)}$
Recall that ${\partial}_{u}{\cal A}=2H(u)\eta(u)$, that ${\partial}_{u}{\cal V}$ is of first order in $u$, and that the leading term in ${\partial}_{u}E$ is $\Delta H(u)$. Using integration by parts we find that the only term involving third order derivatives is $|\nabla H|^{2}$, which was already discussed. Hence we obtain the following result.
Theorem 2.7.
For $u_{0}\in V^{m}_{\varepsilon}$ there exists a $t_{0}=t_{0}(u_{0})$ and a unique solution of (15),
$(u(\cdot,u_{0}),\Lambda(\cdot,u_{0}))\in(L^{2}((0,t_{0}),H^{4}(\mathcal{C}))\cap H^{1}((0,t_{0}),L^{2}(\mathcal{C}))\cap C([0,t_{0}),V_{\varepsilon}^{s})))\times C([0,t_{0}),{\mathbb{R}}^{2}).$
3 Stability of ${\cal C}_{L}$
We consider the linearization of (5) around ${\cal C}_{L}(r)$. This requires some lengthy computations, but since ${\cal C}_{L}(r)$ has constant mean and Gaussian curvatures (with $K=0$), we can explicitly compute the dispersion relation (Lemma 3.1) and subsequently the stability ranges of and bifurcation points from ${\cal C}_{L}(r)$ (Corollary 3.7).
3.1 Linearization and eigenvalues
Since $G(\cdot,\Lambda):V^{4}_{\varepsilon}\to L^{2}$ is an analytic map, we use (11), (12) and (13) to calculate the Frechét derivative via the Gateaux differential in direction $v$, i.e., ${\partial}_{u}G(0,\Lambda)v=\frac{\,{\rm d}}{\,{\rm d}\varepsilon}G(\varepsilon v,\Lambda)|_{\varepsilon=0}$; recall that the Laplace-Beltrami operator on ${\cal C}_{L}((r)$ is $\Delta_{0}={\partial}_{x}^{2}+\frac{1}{r^{2}}{\partial}_{\varphi}^{2}$ in our coordinate system.
Lemma 3.1.
The linearizations of $H(u),K(u)$, and $\Delta_{u}H(u)$ around ${\cal C}_{L}(r)$, i.e., around $u=0$, are given by
${\partial}_{u}H(0)v$ $=-\frac{1}{2}\left(\Delta_{0}v+\frac{1}{r^{2}}v\right),$ (33)
${\partial}_{u}K(0)v$ $=-r^{-1}{\partial}_{x}^{2}v,\text{ and }$ (34)
${\partial}_{u}\Delta_{0}H(0)v$ $=\Delta{\partial}_{u}H(0)v.$ (35)
Proof.
Using (11) with $\varepsilon v$ and sorting by $\varepsilon$ powers gives
$2H(\varepsilon v)=-\frac{\varepsilon(r^{3}v_{xx}+rv_{\varphi\varphi})+{\cal O}(\varepsilon^{2})}{(r^{2}+\varepsilon 2rv+{\cal O}(\varepsilon^{2}))^{3/2}}+\frac{1}{(r^{2}+\varepsilon 2rv+{\cal O}(\varepsilon^{2}))^{1/2}}.$
Differentiating w.r.t. $\varepsilon$ this gives
$\frac{\,{\rm d}}{\,{\rm d}\varepsilon}2H(\varepsilon v)=-\frac{(r^{3}v_{xx}+rv_{\varphi\varphi})(r^{2}+{\cal O}(\varepsilon))^{3/2}+{\cal O}(\varepsilon)}{(r^{2}+\varepsilon 2rv+{\cal O}(\varepsilon^{2}))^{3}}-\frac{rv+{\cal O}(\varepsilon)}{(r^{2}+{\cal O}(\varepsilon))^{3/2}}.$
The denominators are bounded away from zero, hence evaluating at $\varepsilon=0$ gives (33).
For $K$ this works the same way, with
$\frac{\,{\rm d}}{\,{\rm d}\varepsilon}K(\varepsilon v)=\frac{\,{\rm d}}{\,{\rm d}\varepsilon}\frac{-\varepsilon r^{3}v_{xx}+{\cal O}(\varepsilon^{2})}{(r^{2}+2\varepsilon 2rv+{\cal O}(\varepsilon^{2}))^{2}}=\frac{-r^{3}v_{xx}((r^{2}+\varepsilon 2rv+O(\varepsilon^{2}))^{2}+{\cal O}(\varepsilon)}{(r^{2}+\varepsilon 2rv+{\cal O}(\varepsilon^{2}))^{4}}.$
Again evaluating at $\varepsilon=0$ gives (34).
Finally for (35) it is sufficient to notice that
$\frac{\,{\rm d}}{\,{\rm d}\varepsilon}\Delta_{\varepsilon v}H(\varepsilon v)|_{\varepsilon=0}=\left(\frac{\,{\rm d}}{\,{\rm d}\varepsilon}\Delta_{\varepsilon v}H(0)\right)\bigg{|}_{\varepsilon=0}+\Delta_{0}\left(\frac{\,{\rm d}}{\,{\rm d}\varepsilon}H(\varepsilon v)\bigg{|}_{\varepsilon=0}\right)=\Delta_{0}\left(\frac{\,{\rm d}}{\,{\rm d}\varepsilon}H(\varepsilon v)\bigg{|}_{\varepsilon=0}\right),$
since $H(0)$ is a constant and $\Delta_{\varepsilon v}$ is non degenerate for small perturbations. ∎
Inserting the expressions from Theorem 3.1 we obtain
Lemma 3.2.
$G(\cdot,\Lambda)$ is Frechet differentiable at $u=0$, with differential, using (7),
${\partial}_{u}G(0,\Lambda)v=-\frac{1}{2}\Delta^{2}_{0}v-\left(r\lambda_{2}+\frac{1}{r^{2}}\right)\Delta_{0}v-\left(\frac{\lambda_{2}}{r}+\frac{1}{2r^{4}}\right)v+\left(\frac{1}{r}-2c_{0}\right)r^{-1}{\partial}_{x}^{2}v.$ (36)
Over the unbounded ${\cal C}(r)$, ${\partial}_{u}G(0,\Lambda))$ has a continuous spectrum, but by restricting to the compact cylinder ${\cal C}_{L}(r)$ of length $L$, ${\partial}_{u}G(0,\Lambda))$ has a discrete spectrum.
Lemma 3.3.
The eigenvalue problem
${\partial}_{u}G(0,\Lambda)\psi=\mu\psi$ (37)
has a countable set of solutions $(\mu_{mn},\psi_{mn})\in{\mathbb{R}}\times H^{4}_{\text{per}}({\cal C}_{L}(r))$, $(m,n)\in{\mathbb{Z}}^{2}$,
$\psi_{mn}(x,\varphi)=e^{{\rm i}(k_{m}x+n\varphi)}$, $k_{m}=2m\pi/L$, and $\mu_{mn}=\mu(k_{m},n)$ with (38a)
$\mu(k,n,\Lambda)=-\frac{1}{2}\Big{[}k^{2}{+}\frac{n^{2}}{r^{2}}\Big{]}^{2}+\left(r\lambda_{2}{+}\frac{1}{r^{2}}\right)\Big{[}k^{2}{+}\frac{n^{2}}{r^{2}}\Big{]}-\left(\frac{\lambda_{2}}{r}{+}\frac{1}{2r^{4}}\right)-\left(\frac{1}{r}{-}2c_{0}\right)\frac{k^{2}}{r}.$ (38b)
For all $\Lambda\in{\mathbb{R}}^{2}$ the eigenfunctions $\psi_{0\pm 1}\in\ker({\partial}_{u}G(0,\Lambda))$.
Proof.
Recall that $H^{4}_{\text{per}}({\cal C}_{L}(r))$ has a compact embedding in $L^{2}({\cal C}_{L}(r))$, hence the resolvent of (36) is compact. Using the $L^{2}$ basis of eigenfunctions of $\Delta_{0}$,
$\psi_{mn}(x,\varphi)=e^{{\rm i}(k_{m}x+n\varphi)},\quad\Delta_{0}\psi_{mn}=-\Big{[}k_{m}^{2}+\frac{1}{r^{2}}n^{2}\Big{]}\psi_{mn}(x,\varphi),$
we obtain (38b). Clearly $\mu(0,\pm 1,\Lambda)=0$, hence $\sin(\varphi),\cos(\varphi)\in\ker({\partial}_{u}G(0,\Lambda))$. ∎
Using Lemma 2.1, we write the linearization of (15) as
${\partial}_{U}\mathcal{G}(0,\Lambda)U$ $=\begin{pmatrix}{\partial}_{u}G(0,\Lambda)&{\partial}_{\lambda_{1}}G(0,\Lambda)&{\partial}_{\lambda_{2}}G(0,\Lambda)\\ {\partial}_{u}{\cal A}(0)&{\partial}_{\lambda_{1}}{\cal A}(0)&{\partial}_{\lambda_{2}}{\cal A}(0)\\ {\partial}_{u}{\cal V}(0)&{\partial}_{\lambda_{1}}{\cal V}(0)&{\partial}_{\lambda_{2}}{\cal V}(0)\end{pmatrix}\begin{pmatrix}u\\ \sigma_{1}\\ \sigma_{2}\end{pmatrix}$
$=\begin{pmatrix}{\partial}_{u}G(0,\Lambda)u&-2H(0)\sigma_{1}&-\sigma_{2}\\ -\int_{{\cal C}_{L}(r)}2H(0)u\,{\rm d}S&0&0\\ \int_{{\cal C}_{L}(r)}u\,{\rm d}S&0&0\end{pmatrix}$. (39)
The domain of ${\partial}_{U}{\cal G}(0,\Lambda)$ is $H^{4}_{\text{per}}({\cal C}_{L}(r))\times{\mathbb{R}}^{2}$, equipped with the norm $\|U\|_{{\partial}_{(u,\Lambda)}\mathcal{G}}=\|u\|_{H^{4}}+\|\Sigma\|_{{\mathbb{R}}^{2}}$, and altogether we have the generalized eigenvalue problem
${\partial}_{U}\mathcal{G}(0,\Lambda)\Psi=\mu\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}\Psi,\quad\Psi=(\psi,\sigma_{1},\sigma_{2}),$ (40)
[…]
$\Psi_{mn}=\begin{pmatrix}\psi_{mn}\\ 1\\ -2\end{pmatrix},\,(m,n)\in{\mathbb{Z}}^{2}\backslash\{(0,0)\},\text{ and }\Psi_{00}=\begin{pmatrix}0\\ 1\\ -2\end{pmatrix}$.
Remark 3.4.
(a) The eigenvalues $\mu(0,\pm 1,\lambda)=0$ correspond to translations of ${\cal C}_{L}(r)$ in $y$ and $z$ directions. Each eigenvalue of the form $\mu(k_{m},0,\Lambda)$ and $\mu(0,n,\Lambda)$ has multiplicity two and every “double” eigenvalue $\mu(k_{m},n,\Lambda)$ with $m\neq 0\neq n$ has multiplicity four with eigenfunctions $\psi_{\pm m\pm n}$.
(b) The eigenvalues $\mu(k_{m},n,\Lambda)$ do not explicitly depend on $\lambda_{1}$, but implicitly. Recall that $\lambda_{1}=\frac{1}{2}\left(\frac{1}{2r^{2}}-2r\lambda_{2}-2c_{0}^{2}\right)$, hence the parameters are not independent on the cylinder.
(c) In the following we also use the notation $\Psi_{m,n}$ for the first two components of $\Psi_{m,n}$, which are the eigenvectors for the case of just the area constraint being active, cf. Theorem 2.6 and (41) in §3.2.
(d) The crucial difference of the spectrum of (40) and just the PDE part (37) is that $(\psi_{00},1,-2)$ with $\psi_{00}\equiv 1$ is not an eigenvector of ${\partial}_{u}\mathcal{G}(0,\Lambda)$. $\rfloor$
3.2 Center manifold construction
Based on the local existence Theorem 2.6 and the spectral computations from Lemma 3.1 we construct the center manifold for the flow near ${\cal C}_{L}(r)$ which will be the basis for the subsequent stability and bifurcation results. At ${\cal C}_{L}(r)$, $\det{\cal D}=0$, and we cannot perturb ${\cal C}_{L}(r)$ while preserving the area and the volume. Therefore we start with
$\begin{pmatrix}{\partial}_{t}u\\ 0\end{pmatrix}=\begin{pmatrix}\eta^{-1}(u)G(u,\Lambda)\\ q_{1}(u,\Lambda)\end{pmatrix}=:{\cal G}_{\cal A}(u,\Lambda),$ (41)
where $\Lambda=(\lambda_{1},\lambda_{2})$ with $\lambda_{2}$ a fixed external pressure difference. The ${\mathbb{R}}^{3}$ has the Euclidean symmetry group $E(3)$ of rigid body motions, which can be constructed using two smooth curves $\tau,\varrho:{\mathbb{R}}\rightarrow{\mathbb{R}}^{3}$ with $\tau(0)=\varrho(0)=0$, such that the translated and rotated surface is given by
$\tilde{X}=\{y+\tau(\eta)+y\times\varrho(\eta):y\in X\}.$
However, since we only consider normal variations of $X$, admissible rigid body motions are
${\cal M}X=\{X+u\,N:u=\left\langle\tau(\eta)+y\times\varrho(\eta),N(y)\right\rangle\}.$
The dimension of ${\cal M}X$ is given by the dimension of the tangent space $T_{e}{\cal M}X$ at the identity, hence spanned by $\left\langle\tau^{\prime}(0),N(y)\right\rangle$ and $\left\langle y\times\varrho^{\prime}(0),N(y)\right\rangle$. Since $\tau^{\prime}(0)$ and $\varrho^{\prime}(0)$ are arbitrary vectors,
$T_{e}{\cal M}X=\mathrm{span}\{\left\langle e_{i},N\right\rangle,\left\langle y\times e_{i},N\right\rangle,\,i=1,2,3\}.$ (42)
But for ${\cal C}_{L}(r)$ we have
$T_{e}{\cal M}\mathcal{C}_{r}=\mathrm{span}\{\left\langle e_{2},N\right\rangle,\left\langle e_{3},N\right\rangle\}=\mathrm{span}\{\psi_{01},\psi_{0-1}\}$ (43)
where a straightforward calculation shows that translation in and rotation around $e_{1}$ generate tangential motions to ${\cal C}_{L}(r)$, while rotations around $e_{2}$ and $e_{3}$ violate the periodicity in $x$.
In §3.3 we will see that there are parameter values $\lambda_{2}$ such that the spectrum of the linearization of (41) is contained in the negative half plane, with one zero eigenvalue with multiplicity according to (43). The spectrum on the finite cylinder is discrete, with a spectral gap between the zero and first non zero eigenvalue. Hence the spectrum decomposes into the center spectrum $\mathrm{spec}_{c}$ and the stable spectrum $\mathrm{spec}_{s}$. Using the well-posedness Theorem 2.6 we may apply [Sim95, Theorem 4.1] on the existence of a center manifold under the above conditions. Since (43) shows that the center eigenspace coincides with the tangent space of the manifold of translated cylinders, both manifolds also coincide locally. This gives the following results, where again $y$ and $z$ are the euclidian coordinates orthogonal to the cylinder axis $x$.
Lemma 3.5.
At the trivial solution $(0,\lambda_{2})$ with $\lambda_{1}$ and $\lambda_{2}$ related by (7), there exists a smooth center manifold $\mathcal{M}^{c}$ for (41) of dimension at least two. If the kernel of ${\partial}_{(u,\lambda_{2})}{\cal G}_{\cal A}$ is spanned by $\Psi_{0\pm 1}$ then
${\cal M}^{c}=\text{Shifts of ${\cal C}_{L}$ in $y$ and $z$ direction}.$ (44)
Theorem 3.6.
If $\lambda_{2}\in{\mathbb{R}}$ such that the spectrum of ${\partial}_{(u,\lambda_{2})}{\cal G}_{\cal A}$ is contained in $(-\infty,0]$ and $\ker({\partial}_{(u,\lambda_{2})}{\cal G}_{\cal A})$ is spanned by $\Psi_{0\pm 1}$, then there exists $\delta,\omega,M>0$ such that for all $u_{0}\in V^{m}_{\varepsilon}$ with $\|u_{0}\|_{H^{m}}<\delta$ the solution $U=(u,\lambda_{1})$ of (41) exists globally in time and there exist a unique $U_{\infty}\in{\cal M}^{c}$ such that
$\|U(t,U_{0})-U_{\infty}\|_{H^{m}\times{\mathbb{R}}}\leq Me^{-\omega t}\|U_{0}\|_{H^{m}\times{\mathbb{R}}}.$
3.3 Stability and instability
Following Theorem 3.6, to check for stability of ${\cal C}_{L}(r)$ we seek $\Lambda\in{\mathbb{R}}^{2}$ such that all eigenvalues $\mu(k_{m},n,\Lambda)$ are in $(-\infty,0]$, and the only zero eigenvalues belong to $\Psi_{0\pm 1}$ (translations in $y$ and $z$). From here on, to simplify notation we fix $r=1$, and recall that the results for general $r$ follow from rescaling $r\mapsto 1$, $L\mapsto rL$ and $c_{0}\mapsto c_{0}/r$. Practically, we vary $\lambda_{2}$, with $\lambda_{1}$ given by (7) along the trivial branch, and monitor the signs of eigenvalues. Depending on $c_{0},L$, the main questions are:
  • Are there intervals for $\lambda_{2}$ such that ${\cal C}_{L}$ is linearly stable?
  • Which modes destabilize the cylinder at what $\lambda_{2}$?
In Lemma 3.1, we fixed L>0L{>}0 yielding the discrete allowed wave numbers $k_{m}=2\pi m/L$, $m\in{\mathbb{Z}}$. Analytically it is more convenient to return to continuous $k\in{\mathbb{R}}$. Thus we consider
$\mu(k,n,\lambda_{2})=-\frac{1}{2}(k^{2}+n^{2})^{2}+(\lambda_{2}+1)(k^{2}+n^{2})-(\lambda_{2}+1)-(1-2c_{0})k^{2},$ (45)
where due to the reflection (in $x$) symmetry of ${\partial}_{u}G(0)$ in (36) only even powers of $k$ appear, and hence we may restrict to $k\geq 0$, and similarly to $n\geq 0$. Subsequently we reintroduce $k_{m}$ to formulate Corollary 3.7. See also Fig.2 for illustration of the results.
(a)                (b)                (c)
Refer to caption Refer to caption Refer to caption
(d)                   (e)
Refer to caption Refer to caption Refer to caption   Refer to caption
[…]
For $k=0$ we have
$\mu(0,n,\lambda_{2})=-\frac{n^{4}}{2}+\left(\lambda_{2}+1\right)n^{2}-\left(\lambda_{2}+\frac{1}{2}\right),$ (46)
with $\mu(0,1,\lambda_{2})=0$ for all $\lambda_{2}$ (translational eigenvalues), and ${\partial}_{\lambda_{2}}\mu(0,n,\lambda_{2})=n^{2}-1$ is positive for $n>1$. Furthermore $\mu$ is zero at $\lambda_{2}^{(0,n)}=(1-n^{2})^{-1}(-\frac{n^{4}}{2}+n^{2}-\frac{1}{2})$, which is monotonically increasing in $n$. Hence, under axially constant perturbations ${\cal C}_{L}$ is stable for $\lambda_{2}<\frac{3}{2}$, indicated by the violet line in Fig.2.
Next we consider $k>0$ and $n\neq 0$, in particular $n=1$ which turns out to be the most dangerous $n$ for $k\neq 0$. We have
$\mu(k,1,\lambda_{2})$ $=-\frac{1}{2}((k^{2}+1)^{2}+(\lambda_{2}+1)(k^{2}+1)-\left(\lambda_{2}+\frac{1}{2}\right)-(1-2c_{0})k^{2}$
$=-\frac{1}{2}(k_{1}^{2}+1)^{2}+(\lambda_{2}+2c_{0})k^{2}+\frac{1}{2}=-k^{2}(\frac{1}{2}k^{2}+1-\lambda_{2}-2c_{0}).$
One sees directly that $\mu(k,1,\Lambda)$ is strictly increasing in $\lambda_{2}$ and has a zero at $k>0$ for
$\lambda_{2}=\frac{1}{2}k^{2}+1-2c_{0},$ (47)
indicated by the yellow line in Fig.2. In particular, if $c_{0}>-1/4$, then there exists an explicitly computable $L$ large enough so that the $(k_{1},1)$ coiling/buckling instability, $k_{1}=2\pi/L$, in $\lambda_{2}$ is below the wrinkling instability of ${\cal C}_{L}$.
For $k>0$ and $n=0$ (pearling) we have, with $\kappa:=k^{2}$, i.e., $\tilde{mu}(\kappa,\lambda_{2}):=\mu(k^{2},0,\lambda_{2})$,
$\tilde{mu}(\kappa,\lambda_{2})=-\frac{1}{2}\kappa^{2}+(\lambda_{2}+2c_{0})\kappa-\lambda_{2}-\frac{1}{2},$ (48)
and solving $\mu(\kappa,\lambda_{2})=0$ for $\lambda_{2}=\lambda_{2}(\kappa)$ yields
$\lambda_{2}(\kappa)=\frac{1}{2}\frac{\kappa^{2}-4c_{0}\kappa+1}{\kappa-1}=\frac{1}{2}(\kappa-1)+\frac{(1-2c_{0})\kappa}{\kappa-1}.$ (49)
For $c_{0}\neq\frac{1}{2}$, this has a singularity at $\kappa=1$ with
[…]
For $c_{0}=\frac{1}{2}$ we have $\lambda_{2}(\kappa)=-1/2+\kappa/2$. In any case, ${\partial}_{\lambda_{2}}\tilde{mu}(\kappa,\lambda_{2})=\kappa-1$ such that the unstable region is below $\lambda_{2}(\kappa)$ for $0<\kappa<1$ and above $\lambda_{2}(\kappa)$ for $\kappa>1$. See the blue curves/sets in Fig.2. Finally, searching for extrema of $\lambda_{2}(\kappa)$ via ${\partial}_{\kappa}\lambda_{2}(\kappa)=0$ yields
$\kappa_{\pm}=1\pm\sqrt{2-4c_{0}}.$ (50)
Hence, $\kappa_{\pm}$ exist for $c_{0}<1/2$, and $\kappa_{-}\in(0,1)$ for $c_{0}\in(1/4,1/2)$, and (only) in this case we have a finite wave number $k=\sqrt{\kappa}$ instability of ${\cal C}_{L}$, for sufficiently large $L$.
Again, see Fig.2 (a)–(d) for illustration of these instability curves and selected dispersion relations, while in (e) we show the fully unstable case, which will be of interest for the outlook in Fig.14. In the following Corollary we summarize the above, with maximal wavelength refering to the minimal allowed $k_{1}=2\pi/L$.
Corollary 3.7.
The cylinder ${\cal C}_{L}$ is volume and area preserving stable for every $L>0$ if:
  1. (a)
    $c_{0}\in(1/4,1/2)$ and $\lambda_{2}\in(\lambda_{2}^{\text{pearl}},\lambda_{2}^{\text{cb}})$ where $\lambda_{2}^{\text{pearl}}=\lambda_{2}(\kappa_{-})$ with $\kappa_{-}=1-\sqrt{2-4c_{0}}$ and $\lambda_{2}$ from (49), and where $\lambda_{2}^{\text{cb}}=1-2c_{0}$. In particular, ${\cal C}_{L}$ with large $L$ destabilizes near $\lambda_{2}^{\text{pearl}}$ to a finite wavelength pearling, and near $\lambda_{2}^{\text{cb}}$ to maximal wavelength buckling and coiling.
  2. (b)
    $c_{0}\in[-1/4,1/4]$ and $\lambda_{2}\in(\lambda_{2}^{\text{pearl}},\lambda_{2}^{cb})$, where $\lambda_{2}^{\text{pearl}}=-1/2$ and $\lambda_{2}^{\text{cb}}=1-2c_{0}$. In particular, ${\cal C}_{L}$ with large $L$ destabilizes near $\lambda_{2}=-1/2$ to maximal wavelength pearling, and near $\lambda_{2}^{\text{cb}}$ to maximal wavelength coiling and buckling.
  3. (c)
    $c_{0}\in(-\infty,-1/4]$ and $\lambda_{2}\in(\lambda_{2}^{\text{pearl}},\lambda_{2}^{\text{wri}})$, where $\lambda_{2}^{\text{pearl}}=-1/2$ and $\lambda_{2}^{\text{wri}})=3/2$. In particular, ${\cal C}_{L}$ with large $L$ destabilizes near $\lambda_{2}=-1/2$ to maximal wavelength pearling, and at $\lambda_{2}=3/2$ to wrinkling (homogeneous in $x$).
If $c_{0}>1/2$, then for all $\lambda_{2}\in{\mathbb{R}}$ there exists an $L_{0}>0$ such that ${\cal C}_{L}$ with $L>L_{0}$ is spectrally unstable.
Remark 3.8.
The values $\lambda_{2}^{\text{cb}}=1-2c_{0}$ and $\lambda_{2}^{\text{pearl}}=-1/2$ are the asymptotics for $k_{1}\to 0$, i.e., for the infinitely long cylinder. In the following section we return to some finite $L>0$, i.e., some finite $k_{1}=2\pi/L$, and in the numerics in §5 we use, e.g., $L=10$, hence $k_{1}=\pi/5\approx 0.628$, or $L=30$, hence $k_{1}=\pi/15\approx 0.209$. The corresponding values for $\lambda_{2}^{\text{cb}}$ and $\lambda_{2}^{\text{pearl}}$ (maximum wavelength pearling) can then be read off from the orange and blue lines in Fig.2. For, e.g., $c_{0}=0.48$ and $L=30$, the finite wavelength pearling corresponds to $k_{m}=2\pi m/30\approx 0.838$ for $m=4$. $\rfloor$
4 Bifurcations from ${\cal C}_{L}$
We now compute the branches bifurcating at the instabilities given in Corollary 3.7. In §4.1 we give the general bifurcation results, based on the $O(2)\times O(2)$ symmetry of the problem at ${\cal C}_{L}$, and in §4.2 we derive the associated amplitude equations (AEs) and moreover discuss the stability of the bifurcating branches. [^5] We exclude (leave for future work) the exceptional (co–dimension two) cases of combinations $(L,c_{0})$ such that at some $\lambda_{2}$ two different modes become unstable simultaneously. This can for instance happen for large $L$ and $c_{0}\stackrel{{\scriptstyle>}}{{\approx}}-0.25$, namely mode crossing of wrinkling and coiling/buckling.
4.1 General bifurcation results
As a bifurcation problem, the desired primary bifurcation parameter is either $\cal V$} or $\cal A$}. However, these parameters do not change along the trivial $u=0$ branch, and therefore we first choose $\lambda_{2}$ as the primary bifurcation parameter and only consider (41), which we write as
$\eta(u){\partial}_{t}u$ $=G(u,\Lambda),$ (51a)
$0$ $={\cal A}(u)-{\cal A}_{0}.$ (51b)
[…]
$\Psi_{mn}(x,\varphi)=\begin{pmatrix}A_{mn}e^{{\rm i}(k_{m}x+n\varphi)}+\mathrm{c.c.}\\ 0\end{pmatrix}$. (52)
Here $\mu_{00}$ is not present since it corresponds to changes of radius along the cylinder branch. Solving (2) gives (possible) bifurcation points (for completeness with rr–dependence)
$\lambda_{2}=\Bigg{[}\frac{(1-n^{2})}{r^{2}}-k_{m}^{2}\Bigg{]}^{-1}\bigg{(}-\frac{1}{2}\Big{[}k_{m}^{2}+\frac{n^{2}}{r^{2}}\Big{]}^{2}+\frac{1}{r^{2}}\Big{[}k_{m}^{2}+\frac{n^{2}}{r^{2}}\Big{]}-\left(\frac{1}{r}-2c_{0}\right)\frac{k_{m}^{2}}{r}-\frac{1}{2r^{4}}\bigg{)}.$ (53)
Due to symmetries the bifurcation points have multiplicities at least two, and hence the equivariant branching lemma [CL00, GS02, Hoy06] gives the needed framework for the bifurcation analysis.
Theorem 4.1.
At (53) in the system (51) we have equivariant pitchfork bifurcations of steady state solutions $X(\varepsilon)={\cal C}_{L}-\varepsilon u(\varepsilon)N$, $\Lambda(\varepsilon)=\Lambda_{0}+{\cal O}(\varepsilon^{2})$, for $0<|\varepsilon|<\varepsilon_{0}$ sufficiently small. The maps $\varepsilon\mapsto(u(\varepsilon),\Lambda(\varepsilon))\in H^{4}\times{\mathbb{R}}^{2}$, or, equivalently $\varepsilon\mapsto(X(\varepsilon),\Lambda(\varepsilon))\in H^{4}_{\text{per}}({\cal C}_{L},{\mathbb{R}}^{3})\times{\mathbb{R}}^{2}$ are analytic, and we characterize the bifurcations as follows:
  1. (a)
    $n=0$ (pearling):
    $u=\left(A_{m}e^{{\rm i}k_{m}x}+\mathrm{c.c.}\right)+{\cal O}(\varepsilon)$ (54)
    with amplitude $A_{m}\in{\mathbb{C}}$; the isotropy class can be represented by the isotropy subgroup $\Sigma:=\{\sigma\in\Gamma:\sigma X=X\}=\{(0,m_{x},\tau_{\phi},m_{\phi})\}\leqslant O(2)\times O(2)$, where $m_{x}$ stands for reflections in $x$ at $x_{0}$ with maximal or minimal $r(x_{0})$, $\tau_{\phi}$ stands for translations in $\phi$ (rotations), and $m_{\phi}$ for reflections in $\phi$. The fixed point subspace ${\rm Fix}\ \Sigma$ consists of surfaces of revolution.
  2. (b)
    $m=0$ (wrinkling):
    $u=\left(A_{n}e^{{\rm i}n\varphi}+\mathrm{c.c.}\right)+{\cal O}(\varepsilon)$ (55)
    with amplitude $A_{n}\in{\mathbb{C}}$; an isotropy subgroup is $\Sigma=\{(\tau_{x},m_{x},0,m_{\phi})\}$ (translations and reflections in $x$, and reflections in $\phi$). ${\rm Fix}\ \Sigma$ consists of $x$-independent surfaces.
  3. (c)
    $m\neq 0\neq n$ (coiling and buckling): Two types of branches bifurcate. Coiling,
    $u=\left(A_{mn}e^{{\rm i}(k_{m}x+n\varphi)}+\mathrm{c.c.}\right)+{\cal O}(\varepsilon),$ (56)
    with $\Sigma=Z_{2}$ generated by joint reflections in $x$ and $\varphi$ (at suitable $x_{0},\varphi_{0})$, and continuous symmetry $\tau_{\phi,-k_{m}\xi/n}\circ\tau_{x,\xi}$, i.e., spatial translation followed by angular rotation. In particular, spatial translation and angular rotation yield the same non–trivial group orbits.
    The other branch corresponds to buckling as equal amplitude superpositions of coiling,
    $u=\left(A_{mn}e^{{\rm i}(k_{m}x+n\varphi)}+A_{m-n}e^{{\rm i}(k_{m}x-n\varphi)}+\mathrm{c.c.}\right)+{\cal O}(\varepsilon),$ (57)
    with $\Sigma=Z_{2}\times Z_{2}$ (independent reflections in $x$ or $\varphi$ at suitable $x_{0},\varphi_{0}$), and no continous symmetries left.
Proof.
The proof proceeds via the equivariant branching lemma; see, for example, [CL00, Theorem 2.3.2]. By (29), $B(u)$ is elliptic; hence the linearization of $G(u,\Lambda)$ is elliptic as well, and thus the upper left $2\times 2$ block of $\partial_{U}{\cal G}$ is a Fredholm operator. At every $\lambda_{2}$ from (53), the zero eigenvalue is of finite multiplicity, namely multiplicity 2 in (a,b), and multiplicity 4 in (c), with the symmetries acting on the amplitudes $(A_{mn},A_{m-n})\in{\mathbb{C}}^{2}$. For (c), the joint reflection gives the condition that either $A_{mn}$ or $A_{m-n}$ must be zero (coiling), while enforcing a $Z_{2}\times Z_{2}$ symmetry yields $A_{mn}=A_{m-n}$ (buckling). In all cases, the isotropy subgroups acting on the amplitudes yield one-dimensional ${\rm Fix}(\Sigma)$. ∎
Remark 4.2.
The equivariant group actions correspond to the rigid body motions stated in (42). As the solution $X$ on the bifurcating branch has the same symmetry group as ${\cal C}-uN$, with $u$ one of the functions from Theorem 57, we shall compute that, for example, in (a), the translations in $x$ are nontrivial rigid body motions of $X-(\varepsilon A_{m}e^{ik_{m}x}+\mathrm{c.c.}+{\cal O}(\varepsilon^{2}))N$.
Assume that $A_{-m}=A_{m}$ with $A_{m}\in{\mathbb{R}}$, so that $u=\cos(k_{m}x)$. The (not normalized) normal vector is
$\nu=\partial_{x}X(\varepsilon)\times\partial_{\varphi}X(\varepsilon)$ $=\begin{pmatrix}1\\ -\varepsilon k_{m}\sin(k_{m}x)\cos(\varphi)\\ -\varepsilon k_{m}\sin(k_{m}x)\sin(\varphi)\end{pmatrix}\times\begin{pmatrix}0\\ -(1+\varepsilon\cos(k_{m}x))\sin(\varphi)\\ (1+\varepsilon\cos(k_{m}x))\cos(\varphi)\end{pmatrix}+{\cal O}(\varepsilon^{2})$
$=-(1+\varepsilon\cos(k_{m}x))\begin{pmatrix}\varepsilon k_{m}\sin(k_{m}x)\\ \cos(\varphi)\\ \sin(\varphi)\end{pmatrix}+{\cal O}(\varepsilon^{2}).$
Hence $\langle\nu,e_{j}\rangle\neq 0$ for $j=1,2,3$; thus, ${\rm span}\{\langle\nu,e_{j}\rangle,j=1,2,3\}$ is a subspace of the tangent space of the center manifold of $X$. However, as this coincides with the number of equivariant group actions close to the bifurcation point, the tangent space coincides with the three translations. In contrast, a rotation around the $x$-axis yields
$X(\varepsilon)\times e_{1}=\begin{pmatrix}0\\ -(1+\varepsilon\cos(k_{m}x))\sin(\varphi)+{\cal O}(\varepsilon^{2})\\ (1+\varepsilon\cos(k_{m}x))\cos(\varphi)+{\cal O}(\varepsilon^{2})\end{pmatrix}$,
so $\langle X(\varepsilon)\times e_{1},\nu\rangle=0$ and the rigid body motion is trivial. A similar argument applies for (b).
In (c), the 4D kernel generates two different types of branches. For the buckling, a combination of the above arguments yields that translations in $x$ and the rotations in $\varphi$ are nontrivial rigid body motions. For the coiling we have, choosing $u=\cos(k_{m}x+n\varphi)$,
$\partial_{x}X(\varepsilon)$ $=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}-\varepsilon k_{m}\sin(k_{m}x+n\varphi)\begin{pmatrix}0\\ \cos(\varphi)\\ \sin(\varphi)\end{pmatrix}+{\cal O}(\varepsilon^{2}),$
$\partial_{\varphi}X(\varepsilon)$ $=(1+\varepsilon\cos(k_{m}x+n\varphi))\begin{pmatrix}0\\ -\sin(\varphi)\\ \cos(\varphi)\end{pmatrix}-\varepsilon n\sin(k_{m}x+n\varphi)\begin{pmatrix}0\\ \cos(\varphi)\\ \sin(\varphi)\end{pmatrix}+{\cal O}(\varepsilon^{2}),$
so the normal vector is
$\nu=$ $(1+\varepsilon\cos(k_{m}x+n\varphi))\begin{pmatrix}0\\ -\cos(\varphi)\\ -\sin(\varphi)\end{pmatrix}-\varepsilon n\sin(k_{m}x+n\varphi)\begin{pmatrix}0\\ -\sin(\varphi)\\ \cos(\varphi)\end{pmatrix}$
$-\varepsilon k_{m}\sin(k_{m}x+n\varphi)(1+\varepsilon\cos(k_{m}x+n\varphi))\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}+{\cal O}(\varepsilon^{2}).$
This results in a nontrivial translational symmetry in $x$. Similarly,
$X(\varepsilon)\times e_{1}=\begin{pmatrix}0\\ -(1+\cos(k_{m}x+n\varphi))\sin(\varphi)\\ (1+\cos(k_{m}x+n\varphi))\cos(\varphi)\end{pmatrix}+{\cal O}(\varepsilon^{2}),$
such that $\langle X\times e_{1},\nu\rangle=\varepsilon n\sin(k_{m}x+n\varphi)(1+\varepsilon\cos(k_{m}x+n\varphi))+{\cal O}(\varepsilon^{2})$ yields the same action as the translation. These computations are also important for the numerics, as they give the formulas and numbers of phase conditions needed for the numerical continuation in §5. $\rfloor$
4.2 Reduced equation
To further describe the bifurcations from Theorem 4.1, we derive the reduced equations (or amplitude equations AEs) at ${\cal C}_{L}$. These are straight forward computations, but somewhat cumbersome in our geometric setting and thus we partly use Maple. Here we explain the basic steps, and the computations for the pearling and wrinkling instability, but for the coiling vs buckling we relegate some details to Appendix A. Additionally we give the numerical values of the coefficients of the AEs for some of examples, which will later be compared to numerics for the full problem.
According to Theorem 4.1, at “simple” (strictly speaking of multiplicity two, but from here we factor out the translation or rotation) bifurcation points (53), i.e., (m,n)=(m,0)(m,n){=}(m,0) or (m,n)=(0,n)(m,n){=}(0,n), the bifurcating branch is tangential to solutions of $D{\cal G}(0,\Lambda)[(\psi,0)]=0$ where $\cal G$} is the rhs of (51). We thus introduce the small parameter
$\varepsilon=\sqrt{\left|\frac{\lambda_{2}-\lambda_{2}(m,n)}{\beta_{2}}\right|},$ (58)
where $\beta_{2}=\pm 1$ is introduced for later convenience to deal with sub–and supercritical pitchforks, and altogether we use the ansatz
$\varepsilon u=\varepsilon A\psi_{m,n}+\varepsilon^{2}\sum_{(i,j)\in{\cal B}}A_{i,j}\psi_{i,j}+\mathrm{c.c.}+{\cal O}(\varepsilon^{3}),$ (59)
$\lambda_{2}=\lambda_{2}(m,n)+\beta_{2}\varepsilon^{2}+{\cal O}(\varepsilon^{3}),\quad\lambda_{1}=\lambda_{1}(m,n)+\alpha_{2}\varepsilon^{2}+{\cal O}(\varepsilon^{3}),\quad$ (60)
[…]
$\frac{{\rm d}}{{\rm d}T}A=A\left(a\beta_{2}+b|A|^{2}\right).$ (61)
At the “double” BPs $m\neq 0\neq n$ two types of branches bifurcate, and we give the ansatz below.
Remark 4.3.
Near a bifurcation point from a stable branch, the critical mode determines the evolution, while all others are exponentially damped. Hence, in a standard AE setting, the signs of $a$, $\beta_{2}$ and $b$ in (61) determine the stability on the bifurcating branch. Here, from (61) we get stability information for steady states under area constraint, with $\lambda_{2}$ as extrinsic fixed parameter, but this can be extended to stability under area and volume constraints as follows. Close to the bifurcation point $\lambda_{2}(m,n)$ the solution reads $X(\varepsilon)={\cal C}_{L}-\varepsilon A\psi N$. Then the solution is linearly stable iff ${\partial}_{U}{\cal G}(X(\varepsilon),\Lambda)$ (which includes the $q_{1}$ area constraint) has only negative eigenvalues.
For admissible (area and volume preserving) perturbations $\phi:{\cal C}_{L}\rightarrow{\mathbb{R}}$ of $X(\varepsilon)$, we have the necessary and sufficient conditions
$\int_{X(\varepsilon)}H(\varepsilon A\psi)\phi\,{\rm d}S=0\quad\text{ and }\quad\int_{X(\varepsilon)}\varepsilon A\psi g^{-1/2}(\varepsilon A\psi)\phi\,{\rm d}S=0.$ (62)
For the mean curvature we have $H(\varepsilon A\psi)=\frac{1}{2}-\varepsilon A(\mu(m,n)+1)\psi+{\cal O}(\varepsilon^{2})$, and hence admissible perturbations must be orthogonal to $\psi$. In order to construct an admissible perturbation $\tilde{\phi}$, we take a Fourier mode $\psi_{kl}$ as a leading order term. Then we find that if ${\partial}_{u}{\cal A}(\varepsilon)\psi_{kl}={\cal O}(\varepsilon^{2})$ and ${\partial}_{u}{\cal V}(\varepsilon)\psi_{kl}={\cal O}(\varepsilon^{2})$, we can construct a $\phi=\psi_{kl}+g(\varepsilon)$, with $g={\cal O}(\varepsilon)$. Hence, for a branch described by $X(\varepsilon)={\cal C}_{L}-\varepsilon A\psi_{mn}N+{\cal O}(\varepsilon^{2})$, the only possible perturbations have a leading term with $(m,n)\neq(k,l)$. Thus, near bifurcation points associated with loss of stability of ${\cal C}_{L}$ we have the following results:
  • Branches bifurcating from a simple BP (pearling or wrinkling) are stable wrt (5).
  • Solutions on branches from a double BP (buckling and coiling) can only be destabilized by the other kernel vectors. We will see examples of this in §5. $\rfloor$
4.2.1 Pearling instability
For the pearling instability (with for notational simplicity $m=1$) we have
$\varepsilon\,u=\varepsilon A\,\psi_{10}+\varepsilon^{2}\left(\frac{1}{2}A_{0}\psi_{00}+A_{2}\,\psi_{20}\right)+\mathrm{c.c.}+{\cal O}(\varepsilon^{3}),\quad\varepsilon=\sqrt{\left|\frac{\lambda_{2}(1,0)-\lambda_{2}}{\beta_{2}}\right|},$
with $A=A(T),\ T=\varepsilon^{2}t$. Plugging into (51) and sorting by powers of $\varepsilon$, with the abbreviations $\lambda_{2}=\lambda_{2}(1,0)$ and $k=k_{1}$, we obtain the following:
  • $\varepsilon\psi_{10}:$
    $0=-\frac{1}{2}\left(\left(k^{4}+\left(-4c_{0}-2\lambda_{2}\right)k^{2}+2\lambda_{2}+1\right)\right)A.$
  • $\varepsilon^{2}\psi_{00}:$
    $0=\left(-\lambda_{2}-\frac{1}{2}\right)A_{0}+\frac{1}{2}\left(\left(k^{4}+\left(-8c_{0}-2\lambda_{2}-1\right)k^{2}+4\lambda_{2}+5\right){|A|}^{2}\right)-\beta_{2}-\alpha_{2}.$
  • $\varepsilon^{2}\psi_{20}:$
    $0=\left(-\frac{7k^{4}}{4}{+}\frac{\left(-3-8c_{0}+2\lambda_{2}\right)k^{2}}{4}{+}\lambda_{2}{+}\frac{5}{4}\right)A^{2}+\left(-8k^{4}{+}\frac{\left(32c_{0}{+}16\lambda_{2}\right)k^{2}}{4}{-}\lambda_{2}{-}\frac{1}{2}\right)A_{2}$.
  • $\varepsilon^{3}\psi_{10}:$
    ${\hskip-11.38109pt}\frac{{\rm d}}{{\rm d}T}A=$ $\left(\frac{5k^{6}}{2}{+}\frac{\left(1{-}8c_{0}{-}3\lambda_{2}\right)k^{4}}{2}{+}\frac{\left(\frac{7}{2}{+}12c_{0}{+}\lambda_{2}\right)k^{2}}{2}{-}\frac{27}{4}{-}3\lambda_{2}\right)A{|A|}^{2}$
    $+\left(\frac{\left({-}4c_{0}{-}1\right)k^{2}}{2}{+}2\lambda_{2}{+}\frac{5}{2}\right)AA_{0}$
    ${+}\left(4k^{4}{+}4\left({-}\frac{3}{8}{-}\frac{5c_{0}}{2}{-}\frac{\lambda_{2}}{2}\right)k^{2}{+}4\left(\frac{5}{8}{+}\frac{\lambda_{2}}{2}\right)\right)\overline{A}A_{2}+\left(-k^{2}+1\right)\alpha_{2}A.$
From the area constraint we get the condition $0=A_{0}+k^{2}|A|^{2}$ for the $\psi_{00}$ mode, which gives
$\alpha_{2}=\frac{1}{2}\left(k^{4}-8k^{2}c_{0}+4\lambda_{2}+5\right){|A|}^{2}-\beta_{2}.$
We used directly the $\varepsilon$ scale of $\lambda_{2}$, hence we only get back the zero eigenvalue at $\varepsilon\psi_{1}$. The above equations at $\varepsilon^{2}\psi_{00}$ and $\varepsilon^{2}\psi_{20}$ give
$A_{0}$ $=\frac{1}{2\lambda_{2}+1}\left(\left(k^{4}+\left(-8c_{0}-2\lambda_{2}-1\right)k^{2}+4\lambda_{2}+5\right){|A|}^{2}+-2\alpha_{2}-2\beta_{2}\right),\text{ and}$ (63)
$A_{2}$ $=-\frac{A^{2}\left(7k^{4}+\left(3+8c_{0}-2\lambda_{2}\right)k^{2}-4\lambda_{2}-5\right)}{2\left(16k^{4}+\left(-16c_{0}-8\lambda_{2}\right)k^{2}+2\lambda_{2}+1\right).}$ (64)
The dependence on $\beta_{2}$ here comes from the dependence of $\lambda_{1}$ on $\lambda_{2}$. Combining everything leads at $\varepsilon^{3}\psi_{10}$ to the solvability condition (61) with
$a=$ $k^{2}-1,$
$b=$ $\frac{5}{\left(k^{2}{-}1\right)k^{2}\left(4k^{4}{-}5k^{2}{+}4c_{0}{-}1\right)}\Bigg{(}\frac{3}{20}k^{14}{+}\left(4c_{0}{-}\frac{73}{20}\right)k^{12}{+}\left(\frac{8c_{0}}{5}{-}\frac{13}{10}\right)k^{10}$
${+}\left(\frac{32}{5}c_{0}^{2}{-}\frac{52}{5}c_{0}{+}\frac{73}{20}\right)k^{8}{+}\left(\frac{16}{5}c_{0}^{2}{-}8c_{0}{+}\frac{31}{5}\right)k^{6}{+}\left(\frac{4c_{0}}{5}{-}\frac{43}{20}\right)k^{4}{+}\left(\frac{12c_{0}}{5}{-}\frac{3}{2}\right)k^{2}\Bigg{)}.$
Recall that here $k=k_{m}=2m\pi/L$, hence $a<0$ for $L>2m\pi$, and similarly $b<0$ for non small $L$. For instance, $a\approx-0.6$ and $b\approx-3$ in Fig.5, where we compare the AE predictions for $m=1$ pearling with $c_{0}=0$ and $L=10$ with numerics. In any case for $ab<0$ we set $\beta_{2}=-1$ (subcritical case), and otherwise $\beta_{2}=1$ (supercritical case), giving the amplitude $A=\sqrt{\left|\frac{a}{b}\right|}$, with also $\alpha_{2}$ uniquely determined. The change of the volume can be approximated by inserting the amplitude Ansatz, giving $0=A_{0}+A^{2}+\beta_{2}$ at $\varepsilon^{2}\psi_{00}$, and comparison with (63) shows that preserving the area and the volume is not possible along the pearling branch.
4.2.2 Wrinkling instability
For the primary wrinkling instability at $\lambda_{2}(0,2)=3/2$ we have the ansatz
$\varepsilon\,u(\varphi)=\varepsilon Ae^{2{\rm i}\varphi}+\varepsilon^{2}\left(\frac{1}{2}A_{0}+A_{4}e^{4{\rm i}\varphi}\right)+\mathrm{c.c.}+{\cal O}(\varepsilon^{3}),$
again with $\varepsilon=\sqrt{\frac{|\lambda_{2}-\lambda_{2}^{(0,2)}|}{\beta_{2}}}$ and $A=A(T)$, $T=\varepsilon^{2}t$, and obtain the following.
  • $\varepsilon\psi_{02}$:
    $0=3A\left(\lambda_{2}-\frac{3}{2}\right)$.
  • $\varepsilon^{2}\psi_{00}$:
    $0=\frac{1}{4}\left(-40\lambda_{2}+82\right){|A|}^{2}-\left(\lambda_{2}+\frac{1}{2}\right)A_{0}-\beta_{2}-\alpha_{2}$.
  • $\varepsilon^{2}\psi_{04}$:
    $0=\left(-9\lambda_{2}+\frac{369}{4}\right)A^{2}+\left(15\lambda_{2}-\frac{225}{2}\right)A_{4}$.
  • $\varepsilon^{3}\psi_{02}$:
    $\frac{{\rm d}}{{\rm d}T}A=-6\left(\lambda_{2}-\frac{11}{4}\right)AA_{0}+\frac{1}{4}\left(-120\lambda_{2}+1170\right)\overline{A}A_{4}-3A\alpha_{2}-3\left(-\lambda_{2}+\frac{221}{4}\right){|A|}^{2}A$.
From the area constraint at $\varepsilon^{2}\psi_{00}$ we have $0=8|A|^{2}+2A_{0}$, leading to $\alpha_{2}=\frac{3\left(15-4\lambda_{2}\right){|A|}^{2}}{2}-\beta_{2}$. Thus $\lambda^{(0,2)}_{2}=\frac{3}{2}$, $A_{0}=\frac{11}{4}|A|^{2}-\frac{\beta_{2}}{2}-\frac{\alpha_{2}}{2}$, $A_{4}=\frac{7}{8}A^{2}$ and finally $\alpha_{2}=\frac{27}{2}|A|^{2}-\beta_{2}$. This ultimately leads to the solvability condition
$\frac{{\rm d}}{{\rm d}T}A=(3\beta_{2}-15.19|A|^{2})A,$ (65)
which is notably independent of $c_{0}$ and $L$, and yields supercritical wrinkling bifurcations for $\beta_{2}=1$.
4.2.3 Coiling/buckling instability
For the instability with critical wave vector $(m,n)=(1,1)$ and (m,n)=(1,1)(m,n)=(1,-1) we choose the ansatz
$\varepsilon\,u(x,\varphi)=\varepsilon\left(A\psi_{11}{+}B\psi_{1-1}\right)+\varepsilon^{2}\left(\frac{1}{2}A_{00}{+}A_{22}\psi_{22}{+}B_{22}\psi_{2-2}{+}A_{20}\psi_{20}{+}A_{02}\psi_{02}\right)+\mathrm{c.c.}+{\cal O}(\varepsilon^{3}),$ (66)
with $\varepsilon=\sqrt{\frac{|\lambda_{2}-\lambda_{2}^{mm}|}{\beta_{2}}}$ and $T=\varepsilon^{2}t$. Using the amplitude formalism, this gives a cubic system of reduced equations for $A,B$, which by $O(2)\times O(2)$ symmetry must read
$\frac{{\rm d}}{{\rm d}T}A$ $=A(a\beta_{2}+b_{1}|A|^{2}+b_{2}|B|^{2}),$ (67a)
$\frac{{\rm d}}{{\rm d}T}B$ $=B(a\beta_{2}+b_{2}|A|^{2}+b_{1}|B|^{2}).$ (67b)
From that we can directly recover the two types of solutions: Coiling with $A=\sqrt{\left|\frac{a}{b_{1}}\right|}$ and $B=0$ or the other way round; buckling with $A=B=\pm\sqrt{\left|\frac{a}{b_{1}+b_{2}}\right|}$. Here again $\beta_{2}=\pm 1$, depending on the signs of $a$, $b_{1}$, and $b_{1}+b_{2}$.
We have $a=k_{m}^{2}$, but the computation of the coefficients $b_{1}$ and $b_{2}$ naturally is more cumbersome than for the scalar case (pearling and wrinkling), and therefore relegated to Appendix A. Here we instead briefly comment on the bifurcation directions, and stability of the bifurcating branches under volume and area preserving flows, see also Fig.3. The AEs are the same as in the classical $D_{4}$ symmetry problem [Hoy06, §4.3.1] and lead to the same direction of bifurcations. However the stability information gained is completely different, due to the constraints, cf. Remark 4.2. Namely, the Jacobian of the right hand side of (67) at a steady state $(A,B)$ is
$\begin{pmatrix}a+3b_{1}A^{2}+b_{2}B^{2}&2b_{2}AB\\ 2b_{2}AB&a+b_{2}A^{2}+3b_{1}B^{2}\end{pmatrix}$.
For $a>0$, on the coiling solution branch the eigenpairs are
$\sigma_{1}=(a\beta_{2}+3{\rm sign}(b_{1})|a\beta_{2}|)$, $v_{1}=(1,0)^{T}$, and $\sigma_{2}=(a\beta_{2}+\frac{b_{2}}{|b_{1}|}|a\beta_{2}|)$, $v_{2}=(0,1)^{T}$. (68)
The first direction $v_{1}$ is excluded by the constraints as it points along the coiling branch, and hence the coiling is stable for $\frac{b_{2}}{|b_{1}|}<-{\rm sign}(\beta_{2})$. For the buckling the Jacobian has the eigenpairs
$\sigma_{1}=2\frac{b_{2}+b_{1}}{|b_{1}+b_{2}|},v_{1}=(1,1)^{T}$, and $\sigma_{2}=\frac{b_{2}-b_{1}}{|b_{1}+b_{2}|},v_{2}=(1,-1)^{T}$. (69)
From the constraints again $v_{1}$ is excluded and hence the relevant eigenvalue is $\sigma_{2}$, which is negative for $b_{2}<b_{1}$. In §5 we give two comparisons of the predictions from (67) with numerics, namely $c_{0}=0$ and $L=10$ yielding $(a,b_{1},b_{2})\approx(0.39,-0.67,-0.82)$ in Fig.4, and $c_{0}=0.48$ and $L=10$ yielding $(a,b_{1},b_{2})=(0.39,-0.083,0.56)$ in Fig.10. In both cases we get the correct branching and stability predictions, and also very good quantitative agreement.
$b_{1}$$b_{2}$$|A||A||A||A||A||A||A||A|$
Figure 3: Schematic linearized stability diagram of the amplitude equation depending on $b_{1}$ and $b_{2}$. Orange stands for the coiling and green for the buckling branch. Thick lines indicate stable solutions.
5 Numerical examples
We now use numerical continuation and bifurcation methods to corroborate the local results from §4, and to extend them to a more global picture. This includes some secondary bifurcations with interesting stable shapes away from the straight cylinder. The numerical setup follows the stability analysis in subsection 3.3 and §4.2. Namely, we fix a period $L$ (e.g., $L=10$, $L=15$ or $L=30$, see also Remark 3.3), and the area ${\cal A}={\cal A}(X)$, and initially use $\lambda_{2}$ as the primary bifurcation parameter and $\lambda_{1}$ and ${\cal V}(X)$ as secondary (dependent) parameters. The linearized stability of the solutions, however, is then computed for fixed ${\cal A}(X)$ and ${\cal V}(X)$ while $\lambda_{1}$ and $\lambda_{2}$ are free parameters. Due to the radius-length scaling invariance of the Helfrich energy (1) and the area and volume constraint in the dynamical problem, it is useful to introduce the reduced volume
$v={\cal V}(X)/{\cal V}_{0},$ (70)
where ${\cal V}_{0}$ is the volume of the straight cylinder with the same ${\cal A}(X)$, compare [JSL93], and we normalize the energy
$\tilde{E}(X)=\frac{1}{{\cal A}}\int_{X}(H-c_{0})^{2}\,{\rm d}S.$ (71)
Also for our numerics we focus on the three exemplary cases
(a) $c_{0}=0$, (b) $c_{0}=-1$,  and (c) $c_{0}=0.48$
from Fig.2, and mostly we use $L=10$. For (a) and (b) this is no significant restriction as in these cases the primary bifurcations (to pearling, buckling and coiling in (a), and to pearling and wrinkling in (b)) are of long wave type for any (sufficiently large) $L$, i.e., a different $L$ gives the same types of primary bifurcations with different axial (maximal) periods $L$ (and the wrinkling is independent of $L$). However, for (c) and sufficiently large $L$ (preferably an integer multiple of the critical period $\ell\approx 7.5$) we have a finite wave number pearling instability. Thus, besides the case L=10L{=}10 (for comparison with (a) and (b)), for (c) we shall consider different multiples of 7.5 for $L$, namely L=15,L=30L{=}15,L{=}30, and L=45L{=}45, allowing 2, 4, and 6, pearls on the primary branch. The influence of $L$ will then be the behavior of this branch away from bifurcation, which loses stability in a bifurcation to a single pearl branch, and this BP moves closer to the primary BP with increasing $L$. This is further discussed in §6. Additionally, in §5.4 we give an outlook on case (d) $c_{0}=0.75$, from Fig.2, where ${\cal C}_{L}$ is unstable for all $\Lambda$.
In the numerics we (as always) have to find a compromise between speed and accuracy. For the default $L=10$ we start with an initial discretization of the straight cylinder ${\cal C}_{L}$ by $n_{p}\approx 3000$ mesh points; the continuation of the nontrivial branches then often requires mesh refinement (and coarsening), typically leading to meshes of 4000 to 5000 points. For the longer cylinders for $c_{0}=0.48$ we increase this to up to 12000 mesh points. A simple measure of numerical resolution is the comparison of the $\lambda_{2}$ values of the numerical BPs from ${\cal C}_{L}$ with their analytical values from §3.1, and here we note that in all cases we have agreement of at least 4 digits. Moreover, we also compare some numerical nontrivial branches with their descriptions by the AEs, and find excellent agreement.
Remark 5.1.
Some cylindrical solutions have similarities with axisymmetric closed vesicle solutions of (4). We recall from [JSL93] or [MU24b], that from the sphere the first bifurcation is to prolates (rods) and oblates (wafers). For mildly negative or positive $c_{0}$ the prolates are stable near bifurcation and the oblates unstable, and continuing the prolates they become dumbbells (two balls connected by a neck), from which pea shaped vesicles bifurcate which become stable as one of the ball shrinks. For sufficiently negative $c_{0}$, the oblates are stable near bifurcation, and continuing the branch they become discocytes of biconcave red blood cell shape. Here we see some reminiscent behavior: For sufficiently negative $c_{0}$ we find secondary bifurcations of stable wrinkles with “embedded biconcave shapes”, see, e.g., Fig.7E. For sufficiently positive $c_{0}$, the primary pearling bifurcation leads to “strings of dumbbells” rather than strings of (round) pearls, see, e.g., samples C in Figures 11 and 13, or the samples A1–A3 in Fig.14. $\rfloor$
We complement our results by some (area and volume conserving) numerical flows close to steady states, and for these perturb steady states $X_{0}$ in two different ways: one is
$\tilde{X}_{0}=X_{0}+\delta u_{0}N_{0},\quad u_{0}=\frac{1}{1+\xi x^{2}}\cos(\varphi),$ (72)
where $x$ is the coordinate along the cylinder axis, and $\varphi$ the polar angle orthogonal to it, and $\delta\in{\mathbb{R}}$ and $\xi>0$ are parameters, with typically $\xi=1$ and $\delta=-0.125$ or $\delta=-0.25$ (recall that $N$ is the inner normal). The other way is
$\tilde{X}_{0}=X_{0}+\delta u_{0}N_{0},\quad\quad\text{with $u_{0}=\psi_{\text{crit}}$},$ (73)
where $\psi_{\text{crit}}$ is (the $u$ component of) the eigenvector to the largest eigenvalue of the Jacobian at $X_{0}$, and typically $\delta=-0.5$. Due to the negative $\delta$ combined with the inner normal $N$, perturbations of type (72) in general induce a small increase of the initial volume ${\cal V}_{0}={\cal V}+{\cal V}_{\delta}$ and area ${\cal A}_{0}={\cal A}+{\cal A}_{\delta}$, which are conserved under the flow. On the other hand, perturbations of type (73) are area and volume preserving to linear order.
Our numerical flows $t\mapsto(X(t),\Lambda(t))=\Psi_{\text{num}}(t;h)X_{0}$ are ad hoc approximations with a time stepsize $h$ and tolerance tol (see Appendix B), and we summarize the results from our flow experiments as follows:
  1. 1.
    Linearly stable $X_{0}$ (at $({\cal V},{\cal A})$) are also stable under the flow at the perturbed $({\cal V}_{0},{\cal A}_{0})$, and the solution flows back to the respective type (pearling, coiling, …) at $({\cal V}_{0},{\cal A}_{0})$, with a small change in $(\lambda_{1},\lambda_{2})$. See, e.g., Fig.6(a,b).
  2. 2.
    If $X_{0}$ at $({\cal V},{\cal A})$ is linearly unstable, then we generally expect $X(t)$ to flow to the nearest stable steady state (of in general a different class), see Fig.6(c2). However, if we have a significant perturbation of $({\cal A}_{0},{\cal V}_{0})$, then $X(t)$ may still flow to a similar steady state close to $X_{0}$ by adapting $(\lambda_{1},\lambda_{2})$, see, e.g., Fig.6($c_{1}$). This underlines the fact that $\lambda_{1,2}$ are dynamical variables.
  3. 3.
    Naturally, we want $E$ to decrease monotonously in $t$, cf.(6), which if violated gives an easy indicator of poor numerics. The monotonous decrease of $E$ often requires rather small stepsizes and tolerances, and a result is that in many examples a single numerical flow to steady state is as expensive as the computation of the full BD of steady states, which we take as a strong argument for our steady state continuation and bifurcation approach.
  4. 4.
    For $c_{0}=0.75$, where ${\cal C}_{L}$ is always unstable and hence no stable primary bifurcations occur, we can numerically flow to stable finite wavelength pearling, which we subsequently continue in parameters, see Fig.14. However, this is only an outlook, and we leave a detailed and systematic study of this for future work.
5.1 $c_{0}=0$
For $L=10$ and $c_{0}=0$, §3.3 gives that the cylinder destabilizes at $\lambda_{2}\approx-0.95$ to pearling $(k_{1},0)$, and at $\lambda_{2}\approx 1.19$ to coiling/buckling $(k_{1},1)$. Figure 4 shows the BDs and samples starting in (a1) with the energy $E$ over the reduced volume $v$.
(a1) (a2)
Refer to caption Refer to caption Refer to caption
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Figure 4: Basic BD for $c_{0}=0$, $L=10$. (a1) Energy $E$ over reduced volume $v$. The family of cylinders $\mathcal{C}_{10}$ parameterized by $\lambda_{1}$ corresponds to the point where all primary branches bifurcate; A=pearls, B=coils, C=buckling, D=wrinkling; E is a secondary bifurcation from D, yielding a mix of buckling and wrinkling. (a2) Relations between $\lambda_{1}$ and $\lambda_{2}$. The cylinders are the black branch, stable for $\lambda_{2}\in(-0.95,1.19)$. Bifurcation of A at $\lambda_{2}\approx-0.95$, bifurcation of B,C at $\lambda_{2}\approx 1.19$, and bifurcation of D at $\lambda_{2}=3/2$. Samples in (b); here the colors indicate $u$ from the current continuation step, with blue for $u<0$ (yellow for $u>0$) indicating that $X$ moves outward (inward) along the branch. The periodic boundaries are shown in red for better visibility.
Here all branches bifurcate from $v=1$, but (a2) shows that they are well separated in the $(\lambda_{1},\lambda_{2})$ plane. As stated in subsubsection 4.2.1, the blue pearling branch starts stable close to the bifurcation point (sample A1), and stays stable up to rather strong deformations of the original cylinder, sample A2. The area constraint prevents the branch from growing to arbitrary volumes, but due to neck development the continuation becomes more difficult for larger $v$. From the mode $(m,n)=(1,1)$ at $\lambda_{2}\approx 1.19$ we find buckling and coiling. The coefficients in (67) are $(a,b_{1},b_{2})=(0.39,-0.4,-0.67)$, and as predicted (see Remark 5.1) the coiling (sample B) is stable close to bifurcation, while the buckling (sample C) is unstable. This behavior continues to small $v$, with no secondary bifurcations.
Additionally we show one wrinkling branch (sample D) as a later bifurcation from the cylinder, and a secondary bifurcation to a wrinkling–buckling mix (sample E). However, these solutions are all linearly unstable in the $\lambda_{1,2}$ range shown, at the given $\cal V$} and $\cal A$}.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: $c_{0}=0$, comparison of numerics and AE predictions (lines with *), (a) pearling, (b) coils (orange) and buckling (green). Left and center: Amplitude over reduced volume $v$, and over $\lambda_{2}$. Right: $\lambda_{1}$ over $\lambda_{2}$.
Remark 5.2.
The amplitude formalism in subsection 4.2 is only expected to be valid in a small neighborhood of the bifurcation points. In Fig. 5 we show comparisons between the numerical computations and the AE predictions. Here the amplitude formalism for the pearling
$u=\varepsilon A\psi_{10}+{\cal O}(\varepsilon^{2})$, $\lambda_{2}=\lambda_{2}(1,0)-\varepsilon^{2}$,  $\lambda_{1}=\lambda_{1}(1,0)+\alpha_{2}\varepsilon^{2}+{\cal O}(\varepsilon^{4})$,
yields $\frac{{\rm d}}{{\rm d}T}A=A(0.61-2.96A^{2})$ and $\alpha_{2}=0.66A^{2}-1$, and that for the wrinkling $u=\varepsilon A\psi_{01}+{\cal O}(\varepsilon^{2})$, $\lambda_{2}=\lambda_{2}(0,1)+\varepsilon^{2}$, $\lambda_{1}=\lambda_{1}(0,1){+}\alpha_{2}\varepsilon^{2}{+}{\cal O}(\varepsilon^{4})$ yields the universal AE (65), i.e., $\frac{{\rm d}}{{\rm d}T}A=A(3{-}15.19A^{2})$ and $\alpha_{2}=\frac{27}{2}A^{2}{-}1$. For both, pearling and wrinkling, we find an excellent match with the numerics. $\rfloor$
In Fig.6 we show numerical flows from the steady states A–C from Fig.4, perturbed according to (72), with parameters given in Table 1. In all cases we integrate until $\|{\partial}_{t}u\|_{\infty}<0.001$. The plots on the left show the initially perturbed $X_{0}$, and $X$ at the end of the DNS colored by the last update $\delta u$, such that $u_{t}\approx\delta u/h$, where $h$ is the current stepsize. The plots on the right show the evolution of $E(t)$ and $\lambda_{1,2}$, and the relative errors in $\cal A$} and $\cal V$} (with the error in $\cal V$} always two orders of magnitude larger than that in $\cal A$}).
(a), from Fig.4A2
Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
(b), from Fig.4B
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(c1), from Fig.4C, perturbation by (72)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(c2), from Fig.4C, perturbation by (73)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Numerical flows from perturbations of states from Fig.4; initial and ’final’ state, and evolution of $E$, $\lambda_{1,2}$ (only $\lambda_{2}$ in (a) and (c2)), and deviations from initial area and volume; very slow evolution in (c2).
The first two flows start near stable steady states $(X_{0},\lambda_{1,0},\lambda_{2,0})$, and as expected flow back to near $X_{0}$ rather quickly, with partly significant changes of $(\lambda_{1},\lambda_{2})$. However, the third flow (c1) is somewhat surprising. State C from Fig.4 is linearly unstable, but the perturbed $X_{0}$ flows to another buckling by adjusting $\lambda_{1,2}$ at the increased area and volume, where this buckling is stable. On the other hand, in (c2) for the (linearly) area and volume preserving initial perturbation into the most unstable direction (which point towards coiling), the behavior is different: The solution now converges to coiling, with a rather small flow throughout; $\lambda_{1}$ hardly changes at all, $\lambda_{2}$ only very slightly, and the convergence is very slow (since the leading eigenvalue of the target coiling is negative but very close to zero). For all flows, we can use the final state as an initial guess for steady state continuation (e.g., in $\lambda_{2}$), and the obtained numerical steady states then are (volume preserving) linearly stable.
Table 1: Spectral data ($n_{\text{unst}}$=number of ($V$-preserving) unstable directions, $\mu_{\text{crit}}$=least stable/most unstable eigenvalue), perturbation parameter $\delta$ ($\xi{=}1$ throughout), and results for DNS in Fig.6; $A_{0}{=}20\pi,V_{0}{=}10\pi$. Initial perturbations for A–C of type (72), and for C of type (73). $(\lambda_{1},\lambda_{2})_{0}$ are the values at the (perturbed) steady state.
IC type A2 (pearling) B (coiling) C (buckling) C (buckling)
$n_{\text{unst}},\mu_{\text{crit}}$ 0, -2.22 0, -0.018 1, 0.012 1, 0.012
$\delta$ -0.125 -0.25 -0.25 0.5
$(A-A_{0})/A_{0}$ 0.0012 0.0003 0.0125 0.0024
$(V-V_{0})/V_{0}$ 0.0011 0.0004 0.02 0.004
$(\lambda_{1},\lambda_{2})_{0}$ (3.734, -3.165) (-0.999, 1.387) (-0.972, 1.29) (-0.972, 1.29)
$(\lambda_{1},\lambda_{2})_{\text{end}}$ (3.701, -3.133) (-0.999, 1.389) (-1.003, 1.367) (-0.965, 1.25)
5.2 $c_{0}=-1$
For (strongly) negative $c_{0}$, the wrinkling branches and their secondary bifurcations become relevant. Fig.7 shows basic results for $c_{0}{=}-1$ and L=10L{=}10. As before, at negative $\lambda_{2}\approx-2.254$ and $\lambda_{1}\approx 1.5$ (cf.Fig.2), $\mathcal{C}_{r}$ loses stability to pearling, but now at positive $\lambda_{2}=3/2$ and $\lambda_{1}=-9/4$ it loses stability to the wrinkling branch (samples D1,2), while the bifurcation to buckling and coiling is at a later $(\lambda_{1},\lambda_{2})\approx(-3.77,3.021)$ and yields initially unstable branches. The wrinkling is initially stable, but at $\lambda_{2}\approx 1.6$ loses stability to “oblates embedded in wrinkling”, see sample E, and Remark 5.
Refer to caption Refer to caption
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Figure 7: Basic BD for $c_{0}=-1$, $L=10$, and samples.
Figure 8 shows some numerical flows from states from Fig.7, focusing on wrinkling perturbed according to (72) (pearling is again stable, also under rather large perturbations). D2 at its given $({\cal A}_{0},{\cal V}_{0})$ is linearly (weakly) unstable, and the perturbation yields a slow evolution to an oblate wrinkling, with a very small change in $E$. In (b) and (c) we perturb a coiling and a buckling, and in both cases flow to wrinkling, and the bottom line of Fig.8 and similar experiments is that (stable) wrinkling and oblate wrinkling have rather large domains of attraction. For completeness, Fig.9 shows a comparison with the AE predictions from (65) for wrinkling, again with very good agreement, cf. Remark 5.1.
(a), from Fig.7D2
Refer to caption Refer to caption Refer to caption Refer to caption
(b), from Fig.7B
Refer to captionRefer to caption Refer to caption Refer to caption
(c), from Fig.7C
Refer to captionRefer to caption Refer to caption Refer to caption
Figure 8: Numerical flows near states from Fig.7; initial perturbations of type (72) with $\delta=-0.25$. Flow to “embedded oblate”, here at the boundary, in (a), and to wrinkling in (b,c).
Refer to caption Refer to caption Refer to caption
Figure 9: Comparison with AE prediction for primary wrinkling.
5.3 $c_{0}=0.48$
We now consider $c_{0}\in(1/4,1/2)$, specifically $c_{0}=0.48$, cf. Fig.2(c). By Corollary 3.7, the cylinder destabilizes in a finite wavelength pearling instability, and in a long wave instability to coiling and buckling. The finite pearling period is $\ell:=\ell_{p}\approx 7.5$, and thus we will be mostly interested in length $L$ as multiples of $7.5$.
However, for comparison with the cases $c_{0}=0$ and $c_{0}=-1$ we first again choose $L=10$, and in Fig.10 give the BD of buckling and coiling. The coefficients for the AE description (67) now are $(a,b_{1},b_{2})=(0.39,-0.083,0.56)$, and in agreement with Fig.3 the buckling is now stable at bifurcation and the coiling is unstable. Moreover, we find interesting secondary bifurcations on both branches, where the buckling loses (coiling gains) stability, and these two BPs are connected by a stable intermediate branch, see sample C. The comparison with the AEs in (c) again shows the agreement close to bifurcation, but naturally the (cubic order) AEs cannot capture the fold of the coiling in $\lambda_{2}$.
(a) (b)
Refer to caption Refer to caption Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: $c_{0}=0.48$, $L=10$. Numerical BDs and samples in (a,b,c), where C shows a stable connection between the stable regimes of coiling and buckling. Comparison to AE predictions in (d).
(a1) (a2)
Refer to caption Refer to caption Refer to caption
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to caption
Figure 11: Basic BD for $c_{0}=0.48$, $L=15$. (a) Energy $E$ over reduced volume $v$, and $\lambda_{1}$ over $\lambda_{2}$. (b) Samples from primary pearling, A1 stable, A2 unstable. (c) Further samples: B is period doubling from A branch, stable; C is the primary $m=1$ pearling bifurcation, unstable; D is a (1,1)–coiling, stable; E is a (1,1)–buckling, unstable; F is a (2,1)–coiling.
Figure 11 shows the BD and samples for L=15L{=}15. To the right of v=1v{=}1 bifurcates the $k_{2}{\approx}0.838$ pearling branch (at $\lambda_{2}{\approx}-0.243$), and is initially stable (sample A1). However, shortly after ($\lambda_{2}{\approx}-0.3756$) there is a secondary bifurcation to a stable branch with double the period, sample B. On the other hand, the primary m=1m{=}1 pearling branch, sample C, (bifurcating at $\lambda_{2}{\approx}-0.421$) is unstable throughout. [^6] At first sight, C may look similar to the $m=2$ pearling A1 and A2, but the difference is that C has one weak and one stronger constriction, which intuitely makes C one period in a string of dumbbells; see also C in Fig. 13, and Remark 5. The (m,n)=(1,1)(m,n){=}(1,1) coiling (sample D) and buckling (sample E) bifurcate at $\lambda_{2}{\approx}0.138$, and again the coiling is stable, and stays stable to rather small $v$, while the buckling is unstable. Wrinkling does not occur in the parameter regime considered; however, the (compared to $L=10$ before) slightly larger $L=15$ lets the pearling and coiling/buckling BPs move closer together, which is why we chose to also present the $(m,n)=(2,1)$ coiling F (unstable). Figure 12 shows the flow from a perturbation of A2 to a B type solution but rather close to the BP. Therefore, the convergence to steady state is again very slow, as near bifurcation the stable B branch has a leading eigenvalue close to 0.
Refer to captionRefer to caption Refer to captionRefer to caption
Figure 12: Sample flow from perturbation of A2 of type (73) with $\delta=0.5$, going to B type solution.
(a) (b)
Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Figure 13: $c_{0}=0.48$, $L=30$. The $m=4$ (period $\ell=7.5$) pearling A now loses stability at BP1 much closer to its primary bifurcation, to a single “localized” pearling branch, sample B. Further branches: C: $m=2$ ($\ell=15$) primary pearling; E: $m=3$ ($\ell=10)$ primary pearling; D: 2ndary pearling from 2nd BP on $m=1$ branch. F: primary $(m,n)=(1,1)$ coiling.
The main point of Fig.11 is that the choice of $c_{0}=0.48$ together with $L=15$ yields the destabilization of the cylinder to the $k_{2}\approx 0.838$ pearling, but these solutions are only stable in a somewhat narrow range (in $v$, say). To further assess this, in Fig.13 we consider the primary pearling with $\ell=7.5$ on twice the domain $L=30$. The result shows that the range of stable $\ell=7.5$ pearling shrinks with increasing $L$, and this continues for further increase of $L$, yielding the following dependence of the reduced volume $v_{s}$ of single pearling BPs from primary $\ell=7.5$ pearling on $L$:
$L$ 15 39 45 60
$v_{s}$ 1.0081 1.0063 1.0046 1.0035
We also ran some numerical flows from perturbations of A in Fig.13, and similarly at larger $L$, with results as in Fig.12, i.e., convergence to the single pearling solutions. However, these DNS are very expensive, partly due to the larger domain and hence larger scale discretization, but more importantly due to the very slow convergence of the flows.
5.4 Outlook: $c_{0}=0.75$
Lastly, following Fig.2(e) we consider $c_{0}>1/2$, specifically $c_{0}=0.75$ for which ${\cal C}_{L}$ is never stable if $L\geq 2\pi$ ($k_{1}\leq 1$, say), and branches as before still bifurcate from ${\cal C}_{L}$, but not at loss of stability of ${\cal C}_{L}$. Hence they can also still be described by AEs as before, but they cannot be stable close to bifurcation, and we can only expect stable “patterns” at $O(1)$ amplitude. This is more difficult than the analysis and numerics so far, and hence here we only give an outlook on some interesting (numerical) results we obtain, again for a rather small length $L=10$ of the cylinder, and leave more systematic and detailed studies to later work.
Figure 14 shows the BD of $k_{1}\approx 0.628$ pearling (A, blue) and $k_{2}\approx 1.257$ pearling (C, light blue), which are both unstable at bifurcation as they must, and moreover are unstable throughout, and addtionally an $O(1)$ amplitude pearling branch (P, red), where P2 is obtained from DNS starting near A1 (panel (c)). In summary, we see a number of new effects which strongly suggest further study, but here we restrict to the following comments.
(a) (b)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14: Selected branches for $c_{0}=0.75$, $L=10$. (a) Energy $E$ over reduced volume $v$, and $\lambda_{1}$ over $\lambda_{2}$. (b) Samples from primary pearling (blue branch in (a)) all unstable. (c) DNS from near $A_{2}$, converging to a stable pearling $P_{2}$ with unequal pearl sizes. (d) Samples from the stable P branch, and sample from (unstable) primary $m=2$ pearling branch (light blue in (a)).
$\bullet$ We stop the continuations shortly after A3, shortly after C, and shortly after P1 as further continuations there require excessive mesh adaptations. We can continue significantly beyond P3, but our main region of interest is near $v=1$. There are bifurcations on the blue branches which break the rotational symmetry, but we restrict to presenting axisymmetric branches.
$\bullet$ The $k_{1}$ pearling branch A shows two folds in $v$, the first at $v_{f,1}\approx 1.045$, the second at $v_{f,2}\approx 0.98$. In particular, between A1 and A2 it goes through $v=1$ again, while keeping the area $\cal A$} constant throughout. This is not a contradiction to the fact that normal variations of ${\cal C}_{L}$ that fix $\cal A$} and $\cal V$} locally do not exist; the passage through $v=1$ here corresponds to a non–small normal variation.
$\bullet$ The folds on the A branch naturally call for fold continuation in $c_{0}$, in particular for decreasing $c_{0}$ as for $c_{0}<1/2$ we do not see folds on the analogous A branches. We believe that they vanish at some $c_{0}>1/2$ as the blue swallowtail shrinks and the folds collide, but this remains to be studied.
$\bullet$ The flow in (c) starts at A1 with a small pertubation in direction of the most unstable eigenvalue; A1 is below the fold, and there is exactly one positive eigenvalue $\mu\approx 0.2$. In any case, the results of flow experiments seem rather robust wrt the choice of the initial conditions (choice of steady state and perturbation), i.e., we always “land” on the P branch, though naturally at different $v$ and $\lambda_{2}$, or analoga of the P branch at different area $\cal A$}. In the particular case shown, the perturbation of A1 yields a relative change of $\cal V$} of order $10^{-3}$, and a relative change of $\cal A$} of order $10^{-4}$. (The initial perturbation linearly conserves $\cal A$} and $\cal V$}, i.e., is tangential to the constraint manifold, but for any finite perturbation we must expect a (nonlinear) change of $\cal A$} and $\cal V$}. However, the change in $\cal A$} is so small that we ignore it and keep the normalization of $E$ in (71) by the original $\cal A$}.)
$\bullet$ The last plots in (c) suggest almost steady states for $t>20$, and we use the solution at $t=20.95$ as an initial guess for Newton’s method for the steady problem. This indeed takes us to the (spectrally) stable steady state P2, and continuing from here to larger/smaller $v$ yields the red branch.
$\bullet$ It is remarkable that the P branch with two pearls of different sizes is stable, that along the branch the pearls change their relative sizes (different from the primary pearling branches, where only the global amplitude changes), and that in fact P2 attracts (perturbations of) A1. However, some finite size effects are involved here, because some stability of the P branch is lost when repeating the experiment from Fig.14 on larger domains, e.g., $L=20$. In this case, initial conditions for numerical flows must be chosen more carefully, namely closer to $v=1$ to reach the analogue of the P branch (4 pearls on $L=20$), and continuation of the P branch then shows a loss of stability (wrt to still axially symmetric solutions) away from $v=1$. We plan to study in particular these effects in a systematic way in future work, including their relation to the transition from A to B in Fig. 13.
$\bullet$ Shape wise (compare P1 and C in (d)) it seems possible that the P branch connects to the C branch in a period–doubling bifurcation, but samples P1 and C are at rather different values of $\lambda_{1,2}$. In our preliminary experiments so far we could not identify to what other branches the P branch connects, and again we leave this for future work.
6 Discussion
Motivated by experimental results as in Fig.1(a,b), we studied the bifurcations from straight cylinders ${\cal C}_{L}$ of lengths $L$ (and wlog radius $r=1$) in the Helfrich model (5). After showing well–posedness we discussed the linearization at ${\cal C}_{L}$ and derived the amplitude equations for the bifurcations of pearling, wrinkling, and coiling and buckling. Subsequently we used numerical continuation and bifurcation methods to extend the local bifurcations to a more global picture.
Our results initially suggest that (5) besides being very interesting mathematically is a possible model for the experiments in Fig.1(a,b). However, when restricting to branches bifurcating at loss of stability of ${\cal C}_{L}$, there also are strong discrepancies. One is that coiling (and buckling, which as far as we know is not clearly documented in experiments) always bifurcate with the maximum wavelength $L$ allowed by the domain. For $c_{0}>1/4$, pearling can occur with a finite wavelength (Figs.1113), but the stability range of finite wavelength pearling shrinks with increasing $L$, probably to zero for $L\to\infty$. This means that the experimental results in Fig.1(a,b) are not described by the primary branches bifurcating at loss of stability of ${\cal C}_{L}$. On the other hand, when further increasing $c_{0}$ beyond $c_{0}{=}1/2$ we find stable finite wavelength $O(1)$ amplitude pearling, but we only gave an outlook on this in §5.4, and this clearly needs to be studied more systematically and in more detail, and for larger $L$.
Mathematically, the bifurcations of localized pearling from finite wave number pearling, with the pertinent BP moving closer to the primary bifurcation for larger domain size $L$, is similar to the bifurcation of localized patterns from periodic patterns in various dissipative systems, e.g., reaction diffusion (RD) systems or Swift–Hohenberg (SH) type equations. See, e.g., [BK07, HK09, BKL+09, Kno15], and [Uec21, Chapters 8 and 9] for further references and numerics using pde2path. In these systems, the localized patterns are usually attributed to subcritical primary bifurcations of periodic patterns, which are hence unstable, but stabilize after a fold at large amplitude. The localized patterns then show some “snaking” behavior with several folds in which the pattern grows, and with alternating stable and unstable segments of this snaking branch. Moreover, in the limit $L\to\infty$, the BP of the snaking branch converges to the primary BP of the periodic patterns.
The data in (76) for BP1 moving closer to $v=1$ for larger $L$ suggest a similar effect, but there are clear differences to the snaking scenario, and many open questions. In our numerics:
  • After the loss of stability of the primary pearling A–branch, it never regains stability; in particular, we cannot find a fold (in $v$, say) for the A–branch.
  • Similarly, we cannot find a fold or other loss of stability of the localized pearling B–branch.
The numerics here are harder for large $L$ than for periodic or localized 1D (or 2D) patterns in, e.g., reaction diffusion systems, and for instance for $L=60$ we only continue the A–branch and the B–branch to $v=1.01$, such that these results are in some sense still local. As mentioned in §5.4, one way to proceed should be fold continuation in $c_{0}$ of the folds that we find on the A branch at larger $c_{0}$.
All our numerics and associated analysis was for finite (and mostly rather small) $L$. For linearizations of (classical) dissipative systems on intervals $(0,L)$ at some steady state, the spectral gap between the critical eigenvalues and the rest of the spectrum shrinks as $L\to\infty$, and over $\mathbb{R}$} the amplitude equations (ODEs) for finite $L$ must be replaced by modulation equations (PDEs) of Ginzburg–Landau type, incorporating the continuous curve of critical eigenvalues. See, e.g., [SU17, Part IV], and the references therein. For the Helfrich cylinders it is not clear how to proceed for $L\to\infty$. For long wave instabilities (buckling, coiling, wrinkling, and pearling except in a small range of positive $c_{0}$ as in Fig.1113), without the volume constraint the critical curve ${\mathbb{R}}\ni k\mapsto\mu(k,n)$ is maximal at wave number $k=0$, and we believe that instead of (5) a better way to incorporate the volume constraint is to consider the $H^{-1}$ gradient flow for (1) with area constraint. This automatically conserves $\cal V$}, and the critical curve is changed to $k\mapsto\tilde{\mu}(k,n)=k^{2}\mu(k,n)$, such that we expect modulation equations with an (additional) conservation law; see, e.g., [MC00, ZS16, Hil20]. This, however, is also beyond the scope of the present paper.
Appendix A Buckling and coiling amplitude system computations
To derive the coefficients $a,b_{1},b_{2}$ for the amplitude system (67) for the mixed mode instability we as usual plug the extended ansatz (66) into (51) and sort wrt powers of $\varepsilon$.
  • $\varepsilon\psi_{11}$:
    $-\frac{1}{2}k^{2}\left(k^{2}-4c_{0}-2\lambda_{2}+2\right)A=0.$
  • $\varepsilon\psi_{1-1}$:
    $-\frac{1}{2}k^{2}\left(k^{2}-4c_{0}-2\lambda_{2}+2\right)B=0$.
  • $\varepsilon^{2}\psi_{00}$:
    $0=\left(-\frac{1}{2}+\frac{k^{4}}{2}+\frac{\left(5-8c_{0}-2\lambda_{2}\right)k^{2}}{2}-\lambda_{2}\right)({|A|}^{2}+{|B|}^{2})-(\lambda_{2}+1/2)A_{00}-\beta_{2}-\alpha_{2}.$
  • $\varepsilon^{2}\psi_{22}$:
    $0=\left(-\frac{7k^{4}}{4}{+}\frac{\left(-8c_{0}{+}2\lambda_{2}{+}19\right)k^{2}}{4}{-}\frac{3\lambda_{2}}{2}{+}\frac{9}{4}\right)A^{2}+\left(-8k^{4}{+}\frac{\left(16\lambda_{2}{+}32c_{0}{-}64\right)k^{2}}{4}{+}3\lambda_{2}{-}\frac{9}{2}\right)A_{22}.$
  • $\varepsilon^{2}\psi_{2-2}$:
    $0=\left(-\frac{7k^{4}}{4}+\frac{\left(-8c_{0}{+}2\lambda_{2}{+}19\right)k^{2}}{4}-\frac{3\lambda_{2}}{2}{+}\frac{9}{4}\right)B^{2}+\left(-8k^{4}{+}\frac{\left(16\lambda_{2}{+}32c_{0}{-}64\right)k^{2}}{4}+3\lambda_{2}{-}\frac{9}{2}\right)B_{22}$.
  • $\varepsilon^{2}\psi_{20}$:
    $0=\left(-\frac{7k^{4}}{2}{+}\frac{\left(2\lambda_{2}+8c_{0}+3\right)k^{2}}{2}{-}\lambda_{2}{-}\frac{1}{2}\right)BA+\left(-8k^{4}+\frac{\left(16c_{0}{+}8\lambda_{2}\right)k^{2}}{2}{-}\lambda_{2}{-}\frac{1}{2}\right)A_{20}$.
  • $\varepsilon^{2}\psi_{02}$:
    $0=4A\left(\frac{k^{4}}{8}+\left(c_{0}-\frac{\lambda_{2}}{4}+\frac{5}{8}\right)k^{2}-\frac{3\lambda_{2}}{4}+\frac{9}{8}\right)\overline{B}+\left(3\lambda_{2}-\frac{9}{2}\right)A_{02}$.
  • $\varepsilon^{3}\psi_{11}$:
    ${\hskip-28.45274pt}\frac{{\rm d}}{{\rm d}T}A=$ $\bigg{[}(-k^{2}\alpha_{2}-4\left(-\frac{5k^{6}}{8}+\left(c_{0}+\frac{3\lambda_{2}}{8}-2\right)k^{4}+\left(-\frac{c_{0}}{2}+\frac{5\lambda_{2}}{8}+\frac{23}{16}\right)k^{2}-\frac{3\lambda_{2}}{4}+\frac{9}{8}\right){|A|}^{2}$
    ${-}8\left({-}\frac{5k^{6}}{8}{+}\left(c_{0}{+}\frac{3\lambda_{2}}{8}\right)k^{4}{+}\left(\frac{3c_{0}}{2}{-}\frac{3\lambda_{2}}{8}{+}\frac{15}{16}\right)k^{2}{-}\frac{3\lambda_{2}}{4}{+}\frac{9}{8}\right){|B|}^{2}{-}4k^{2}\left(\frac{1}{2}c_{0}{-}\frac{3}{8}\right)A_{00}\bigg{]}A$
    ${+}6\left(\left(c_{0}{+}\frac{1}{4}\right)k^{2}{-}\left(\lambda_{2}{-}\frac{3}{2}\right)\right)A_{02}B{-}10\left(-\frac{2k^{4}}{5}{+}\left(c_{0}{+}\frac{\lambda_{2}}{5}{-}\frac{49}{20}\right)k^{2}{+}\frac{3\lambda_{2}}{5}{-}\frac{9}{10}\right)A_{22}\overline{A}$
    $-2k^{2}\left(-2k^{2}+c_{0}+\lambda_{2}-\frac{1}{4}\right)A_{02}\overline{B}.$
  • $\varepsilon^{3}\psi_{1-1}$:
    ${\hskip-28.45274pt}\frac{{\rm d}}{{\rm d}T}B=$ $\bigg{[}{-}k^{2}\alpha_{2}{-}4\left({-}\frac{5k^{6}}{8}{+}\left(c_{0}{+}\frac{3\lambda_{2}}{8}{-}2\right)k^{4}{+}\left(-\frac{c_{0}}{2}{+}\frac{5\lambda_{2}}{8}{+}\frac{23}{16}\right)k^{2}{-}\frac{3\lambda_{2}}{4}{+}\frac{9}{8}\right){|B|}^{2}$
    ${-}8\Big{(}{-}\frac{5k^{6}}{8}{+}\left(c_{0}{+}\frac{3\lambda_{2}}{8}\right)k^{4}{+}\left(\frac{3c_{0}}{2}{-}\frac{3\lambda_{2}}{8}{+}\frac{15}{16}\right)k^{2}{-}\frac{3\lambda_{2}}{4}{+}\frac{9}{8}\Big{)}{|A|}^{2}{-}2k^{2}\left(c_{0}{-}\frac{3}{4}\right)A_{00}\bigg{]}B$
    $+6\left(\left(c_{0}{+}\frac{1}{4}\right)k^{2}{-}\lambda_{2}{+}\frac{3}{2}\right)A\overline{A_{02}}{-}10\bigg{(}{-}\frac{2k^{4}}{5}{+}\left(c_{0}{+}\frac{\lambda_{2}}{5}{-}\frac{49}{20}\right)k^{2}+\frac{3\lambda_{2}}{5}-\frac{9}{10}\bigg{)}B_{22}\overline{B}$
    ${-}2k^{2}\left({-}2k^{2}{+}c_{0}{+}\lambda_{2}{-}\frac{1}{4}\right)A_{20}\overline{A}.$
From the area constraint we get, at $\varepsilon^{2}\psi_{00}$, $0=\left(k^{2}+1\right)(|A|^{2}+|B|^{2})+A_{00}$, and using $\lambda_{2}^{(1,1)}=\frac{1}{2}k^{2}-2c_{0}+1$ we solve the algebraic equations for the $\varepsilon^{2}$ modes:
$A_{00}=-\frac{\left(4k^{2}c_{0}-2k^{2}-4c_{0}+3\right)}{k^{2}-4c_{0}+3}(|A|^{2}+|B|^{2})-\frac{2\alpha_{2}+2\beta_{2}}{k^{2}-4c_{0}+3},$
$A_{22}=-\frac{\left(2k^{4}+4k^{2}c_{0}-6k^{2}-4c_{0}-1\right)}{2\left(4k^{4}+7k^{2}+4c_{0}+1\right)}A^{2},\quad B_{22}=-\frac{\left(2k^{4}+4k^{2}c_{0}-6k^{2}-4c_{0}-1\right)}{2\left(4k^{4}+7k^{2}+4c_{0}+1\right)}B^{2},$
$A_{20}=-\frac{\left(6k^{4}-4k^{2}c_{0}-4k^{2}-4c_{0}+3\right)}{12k^{4}-7k^{2}-4c_{0}+3}AB,\quad A_{02}=-\frac{\left(4k^{2}c_{0}+4c_{0}+1\right)}{k^{2}-4c_{0}-1}A\overline{B}.$
For the Lagrange multiplier expansion we get
$\alpha_{2}=\frac{k^{2}\left(k^{2}-8c_{0}+6\right)}{2}{|A|}^{2}+\frac{k^{2}\left(k^{2}-8c_{0}+6\right)}{2}{|B|}^{2}-\beta_{2},$
and altogether this gives the solvability condition (67) for $A$ and $B$ at $\varepsilon^{3}$ with
$a=$ $k^{2},$
$b_{1}=$ $\frac{5}{4}k^{6}+5k^{4}c_{0}+\frac{3}{4}k^{4}+9k^{2}c_{0}-\frac{33}{4}k^{2}-6c_{0}-\frac{3}{2}$
$+\frac{3\left(2k^{4}+4k^{2}c_{0}-6k^{2}-4c_{0}-1\right)\left(-\frac{1}{2}k^{4}+k^{2}c_{0}-\frac{13}{4}k^{2}-2c_{0}-\frac{1}{2}\right)}{4k^{4}+7k^{2}+4c_{0}+1},$
$b_{2}=$ $3k^{6}+4k^{4}c_{0}-6k^{4}-16k^{2}c_{0}-3k^{2}-12c_{0}-3$
$+\frac{6\left(4k^{2}c_{0}+4c_{0}+1\right)\left(-\left(c_{0}+\frac{1}{4}\right)k^{2}+\frac{k^{2}}{2}-\frac{1}{2}-2c_{0}\right)}{k^{2}-4c_{0}-1}$
$+\frac{2\left(6k^{4}-4k^{2}c_{0}-4k^{2}-4c_{0}+3\right)k^{2}\left(-\frac{3k^{2}}{2}-c_{0}+\frac{3}{4}\right)}{12k^{4}-7k^{2}-4c_{0}+3}.$
Appendix B Numerical algorithms
[…]
$L_{kl}=\int_{X}\nabla\phi_{k}\nabla\phi_{l}\,{\rm d}S=-\frac{1}{2}(\cot\alpha_{kl}+\cot\beta_{kl}).$ (77)
approximates the (positive definite) Laplace–Beltrami operator, while for the FEM mass matrix $M$ we use the so–called Voronoi mass matrix. Then, e.g., $MH(X)=\frac{1}{2}\left\langle LX,N\right\rangle$, where the vertex normals $N$ are obtained from interpolating the normals of the adjacent triangles. For the Gaussian curvature $K$ we use a discrete version of Gauss–Bonnet, and refer to [BGN19] for more sophisticated schemes. See [MU24b, §2.2] for comments on and illustrations of convergence of the used FEM.
Technically, for the 4th order problem (5) we use a mixed formulation as two 2nd order problems, first analyzed in [Rus05] (in a more complicated scheme); see [SGJW22] for a convergence analysis in the frame of the biharmonic equation. Additionally to the PDE $\eta(u){\partial}_{t}u=G(u,\Lambda)$ and the constraints $q(u)=0$, the continuation needs phase conditions (PCs), for which we introduce additional Lagrange multipliers $s_{x},s_{y},s_{z}$ and $s_{rx}$ pertaining to translations in $x,y$ and $z$, respectively, and rotations around the xx–axis. The yy–and zz–translation PCs are always on, the xx–translations is on for pearling and coiling, the xx–rotation is on for the wrinkling, and xx–translation and rotation are on for the buckling. In all continuations, the associated Lagrange multipliers stay ${\cal O}(10^{-6})$. For the discrete systems, an effect of all the constraints (also of $q(u)=0$) is that the linear systems (to be solved in Newton loops and for eigenvalue computations) are bordered, with border widths up to six rows/columns, and we usually use bordered elimination [Uec21, §4.5.1] for their solution.
Mesh handling
The initial cylinder in discretized by a standard mesh from pde2path, mapped via a local map on the cylinder. The number of nodes in $x$ and $\varphi$ direction is chosen to minimize the mesh–quality quotient
$q=\max_{T\in{\cal T}}\frac{e_{M}(e_{1}+e_{2}+e_{3})}{2|T|},$ (78)
with $e_{j}$ the edge lengths, $e_{M}$ the maximal edge length, and $|T|$ the triangle areas. The triangulation is oriented inwards, matching with our analytical convention. On the bifurcating branches triangles distort, due to the change of X. We use mesh adaptation by refinement and coarsening of X and tri based on (78) and triangle areas. Namely, we refine if the area is above a threshold, typically 1.3 times the size of the original cylinder triangles. For the conservation of the periodic boundary, we enforce the same refinement at “opposite” boundary triangles. For the coarsening (78) is the control parameter, activated if $q>8$. We select triangles smaller than a threshold $A_{t}$, typically $A_{t}=\max_{T\in{\cal T}}|T|/3$ and collapse acute triangles along the shortest edge, while obtuse triangles are handled with edge flipping.
Periodicity in $x$
During the continuation of bifurcating branches, it is not necessary that the boundary of $X$ stays in a plane. However, for mesh refinement along the branch the legacy box2per setting looks for periodic boundary nodes in planes. For that reason we only allow an $(y,z)$ displacement at the boundary, by setting $N_{x}=0$.
Extracting the amplitude
For the comparison between the numerics and the AE predictions, we evaluate the deviation at each mesh point. In each continuation step we obtain $X+uN$ as the new solution, but multiple steps result in tangential movement of mesh points relative to the cylinder. Therefore we project each mesh point back onto the cylinder orthogonally to $Cp=\left(p_{1},\frac{p_{2}}{\|(p_{2},p_{3})\|},\frac{p_{3}}{\|(p_{2},p_{3})\|}\right)^{T}.$ From this projection, we can measure the normal deviation from the cylinder and compare it with the amplitude approximation.
On the volume in the numerics
To compute the volume of (not necessarily straight) cylinders $\cal C$} we split them into two domains $\Omega_{1}$ and $\Omega_{2}$, see Fig.15, where $\Omega_{2}=\Omega_{2,a}\cup\Omega_{2,b}$ consists of two equal cones, from $(x,y,z)=(0,0,0)$ to the (in general not circular) left and right disk at $x=\pm L/2$. The idea is to compute $V={\rm vol}(\Omega_{1})+2{\rm vol}(\Omega_{2a})$ by the divergence theorem ${\rm vol}(\Omega)=\frac{1}{3}\int_{\Omega}\mathrm{div}(x)\,{\rm d}x=\frac{1}{3}\int_{{\partial}\Omega}\left\langle x,N\right\rangle\,{\rm d}S$. In the splitting ${\cal C}=\Omega_{1}\cup\Omega_{2}$, we have $\left\langle x,N\right\rangle=0$ along the inner “conical” boundaries ${\partial}{\Omega_{1}}\cap{\partial}{\Omega_{2}}$, such that ${\rm vol}(\Omega_{1})=\frac{1}{3}\int_{X}\left\langle(x,y,z),N\right\rangle\,{\rm d}S$, and ${\rm vol}(\Omega_{2a})=\frac{1}{3}\int_{X_{\text{left}}}\left\langle(-L/2,0,0)^{T},(-1,0,0)^{T}\right\rangle\,{\rm d}(y,z)=\frac{L}{2}|X_{\text{left}}|$, and for the latter we can apply Gauss’ theorem again to obtain
${\rm vol}(\Omega_{2a})=\frac{L}{6}\frac{1}{2}\int_{{\partial}X_{\text{left}}}\left\langle(y,z),(\nu_{1},\nu_{2})\right\rangle\,{\rm d}s=\frac{L}{12}\int_{{\partial}X_{\text{left}}}\left\langle(X_{2},X_{3}),(N_{2},N_{3})/\|(N_{2},N_{3})\|\right\rangle\,{\rm d}s.$ (79)
Thus, altogether,
${\rm vol}({\cal C})=\frac{1}{3}\int_{X}\left\langle(x,y,z),N\right\rangle\,{\rm d}S+\frac{L}{6}\int_{{\partial}X_{\text{left}}}\left\langle(X_{2},X_{3}),(N_{2},N_{3})/\|(N_{2},N_{3})\|\right\rangle\,{\rm d}s.$ (80)
Figure 15: Sketch for volume computation.
In the numerics we also need the derivative of $V({\cal C}+uN)$ wrt $u$. For $\Omega=\Omega_{0}+uN$ and ${\rm vol}(\Omega)=\frac{1}{3}\int_{{\partial}\Omega}\left\langle X,N\right\rangle\,{\rm d}S$ we have ${\partial}_{u}{\rm vol}(\Omega)v=\int_{\Omega}v\,{\rm d}S$, which can be used for ${\partial}_{u}{\rm vol}(\Omega_{1})$, and for ${\partial}_{u}{\rm vol}(\Omega_{2a})$ we proceed as follows. We choose a positive oriented arc length parametrization $\gamma:[0,2\pi)\rightarrow{\partial}X_{\text{left}}$. Then the (unit) tangent and (unit) normal vectors are $T(s)=\begin{pmatrix}\gamma_{1}^{\prime}(s),\gamma_{2}^{\prime}(s)\end{pmatrix}^{T}$ and $\nu(s)=\begin{pmatrix}-\gamma_{2}^{\prime}(s),\gamma_{1}^{\prime}(s)\end{pmatrix}^{T}$. For a perturbation $u$ we consider $\tilde{\gamma}(s)=\gamma(s)+\delta u(s)\nu(s)=\begin{pmatrix}\gamma_{1}(s)-\delta u(s)\gamma^{\prime}_{2}(s)\\ \gamma_{2}(s)+\delta u(s)\gamma^{\prime}_{1}(s)\end{pmatrix}$, with unit normal vector $\tilde{\nu}=\begin{pmatrix}-\tilde{\gamma}^{\prime}_{2}(s)\\ \tilde{\gamma}^{\prime}_{1}(s)\end{pmatrix}/\|\tilde{\gamma}^{\prime}(s)\|$. Then
$\frac{\,{\rm d}}{\,{\rm d}\delta}\bigg{|}_{\delta=0}$ $\int_{0}^{2\pi}\left\langle\tilde{\gamma}(s),\tilde{\nu}(s)\right\rangle\|\tilde{\gamma}^{\prime}(s)\|\,{\rm d}s=\frac{\,{\rm d}}{\,{\rm d}\delta}\bigg{|}_{\delta=0}\int_{0}^{2\pi}\tilde{\gamma}_{2}(s)\tilde{\gamma}^{\prime}_{1}(s)-\tilde{\gamma}_{1}(s)\tilde{\gamma}^{\prime}_{2}(s)\,{\rm d}s$
$=\int_{0}^{2\pi}\frac{\,{\rm d}}{\,{\rm d}\delta}\bigg{|}_{\delta=0}\Big{[}\big{(}\gamma_{2}(s)+\delta u(s)\gamma_{1}^{\prime}(s)\big{)}\Big{(}\gamma_{1}^{\prime}(s)-\delta\big{(}u(s)\gamma_{2}^{\prime}(s)\big{)}^{\prime}\Big{)}$
${\hskip 42.67912pt}-\big{(}\gamma_{1}(s)-\delta u(s)\gamma_{2}^{\prime}(s)\big{)}\Big{(}\gamma_{2}^{\prime}(s)+\delta\big{(}u(s)\gamma_{1}^{\prime}(s)\big{)}^{\prime}\Big{)}\Big{]}\,{\rm d}s$
$=\int_{0}^{2\pi}u(s)\gamma_{1}^{\prime}(s)^{2}-\gamma_{2}(s)(u(s)\gamma_{2}^{\prime}(s))^{\prime}+u(s)\gamma_{2}^{\prime}(s)^{2}-\gamma_{1}(s)(u(s)\gamma_{1}^{\prime}(s))^{\prime}\,{\rm d}s$
$=2\int_{0}^{2\pi}u(s)\,{\rm d}s,$
which can be easily implemented using the (1D) mass matrix of the discretization of ${\partial}X_{\text{left}}$.
Numerical flows
We give a brief description of our ad hoc method for numerical flows (aka direct numerical simulation DNS) for (5), and refer to [BMN05, BGN08, BGN20, BGN21] and the references therein for more sophisticated and provably stable FEM schemes for geometric flows, including the Helfrich flow (5) and extensions, and in particular also dealing with tangential motion of mesh points. Here we only consider normal displacements, and approximate the normal velocity as
$V_{t}=\left\langle{\partial}_{t}X,N\right\rangle\approx\left\langle\frac{1}{h}M_{n}(X_{n+1}{-}X_{n}),N_{n}\right\rangle=\frac{1}{h}\left\langle M_{n}(X_{n}{+}u_{n+1}N_{n}{-}X_{n}),N_{n}\right\rangle=\frac{1}{h}M_{n}u_{n+1},$
where $h$ is the current stepsize, $n$ the time stepping index, and $M_{n}$ the mass matrix. Setting $U_{n}=(u_{n},\Lambda_{n})$, the PDE $V_{t}=G(X(t),\Lambda(t))$, together with the (volume and area) constraints $q(u_{n+1})=0$ (translational and rotational phase conditions play no role in the DNS), yields the nonlinear system
${\cal G}(U_{n+1}):=\begin{pmatrix}\frac{1}{h}M_{n}u_{n+1}+G(U_{n+1})\\ q(u_{n+1})\end{pmatrix}\stackrel{{\scriptstyle!}}{{=}}\begin{pmatrix}0\\ 0\end{pmatrix}\in{\cal H}\times{\mathbb{R}}^{n_{q}}$ (81)
[…]
Additionally we remark that Newton loops for (81) often fail to converge near straight cylinders (or more generally CMC surfaces), where ${\partial}_{u}q_{1}$ and ${\partial}_{u}q_{2}$ become linearly dependent ($\cal D$} from (20) becomes singular), but this is not a major concern for us as we are more interested in the flow towards nontrivial steady states. See, e.g., [BGN08, Scheme (2.8)] for dealing with the constraints when approaching CMC surfaces, which essentially consists in setting $\lambda_{1,n+1}=\lambda_{1,n}$ and only solving for $\lambda_{2,n+1}$ with $q(u_{n+1})=q_{2}(u_{n+1})$. Moreover, the DNS in §5 are all uncritical wrt to mesh quality. However, for stronger initial perturbations we expect that allowing and controlling tangential motion of mesh points (or other ways of mesh–adaptation) will be crucial, see also, e.g., the numerical flows for closed vesicles in [MU24a].